### v2

# On twisted Kawamata’s semi-positivity and finite generation of generalized canonical rings

Abstract.
We give a twisted version of the Kawamata semi-positivity theorem by the $\mathbb{Q}$-line bundle with a vanishing Lelong number at every point. Moreover, we apply the result for the finite generation problem of canonical rings for Birkar’s generalized pairs.
Key words and phrases:
canonical ring, canonical bundle formula, Lelong number
2010 Mathematics Subject Classification:
14D06, 14E30, 32J25
1. Introduction
Throughout this paper, we work over $\mathbb{C}$, the complex number field.
In this paper, we prove the following theorem, which we call ”twisted Kawamata’s semi-positivity theorem”:
Theorem 1.1 (Twisted Kawamata’s semi-positivity theorem).
Let $f:X\to Y$ be a surjective morphism of projective varieties with connected fibers. Suppose $X$ is quasi-smooth and $Y$ is smooth. Let $P=\sum_{j}P_{j}$ and $Q=\sum_{l}Q_{l}$ be reduced effective divisors on $X$ and $Y$, respectively, such that $f^{-1}(Q)\subseteq P$.
Let $M$ be a ${\mathbb{Q}}$-divisor on $X$, which, as a $\mathbb{Q}$-line bundle, admits a singular Hermitian metric with semi-positive curvature and whose local weight functions have vanishing Lelong numbers at any points in $X$ (see 2.2.4 for the definition). In particular, $M$ is nef.
Let $D=\sum_{j}d_{j}P_{j}$ be a $\mathbb{Q}$-divisor ($d_{j}$’s may be negative or zero), which satisfies the following conditions (1)–(4):
  • (1)
    $f:(X,P)\to(Y,Q)$ is a well prepared model ($f:(X,P)\to(Y,Q)$ is toroidal and $f$ is equi-dimensional, in fact $f$ is flat),
  • (2)
    $d_{j}<1$,
  • (3)
    $\dim_{\mathbb{C}(\eta)}f_{*}\mathcal{O}_{X}(\ulcorner-D\urcorner)\otimes_{\mathcal{O}_{Y}}\mathbb{C}(\eta)=1$, where $\eta$ is the generic point of $Y$, and
  • (4)
    $K_{X}+D+M\sim_{\mathbb{Q}}f^{*}(K_{Y}+L)$ for some $\mathbb{Q}$-divisor $L$ on $Y$.
Let (5)
$f^{*}Q_{l}$ $=$ $\sum_{j}w_{lj}P_{j},\ \text{where}\ w_{lj}\in{\mathbb{Z}}_{>0},$
$\bar{d}_{j}$ $=$ $\frac{d_{j}+w_{lj}-1}{w_{lj}}(\in{\mathbb{Q}}_{>1})\ \text{if}\ f(P_{j})=Q_{l},$
$\delta_{l}$ $=$ $\mathrm{max}\{\bar{d_{j}}|f(P_{j})=Q_{l}\},$
$\Delta_{Y}$ $=$ $\sum\delta_{l}Q_{l},\ \text{and}$
$M_{Y}$ $=$ $L-\Delta_{Y}.$
By construction, the pair $(Y,{\Delta}_{Y})$ is sub-klt and $K_{X}+D+M\sim_{\mathbb{Q}}f^{*}(K_{Y}+{\Delta}_{Y}+M_{Y})$. We sometimes call $M_{Y}$ (resp. $\Delta_{Y}$) the moduli part (resp. discriminant part).
Then, the ${\mathbb{Q}}$-divisor $M_{Y}$ admits, as a ${\mathbb{Q}}$-line bundle, a singular Hermitian metric with semi-positive curvature and whose local weight functions have vanishing Lelong numbers at any points in $Y$. In particular, $M_{Y}$ is nef.
The property (3) is, in fact, to say that $f_{*}{\mathcal{O}}_{X}(\ulcorner{-D}\urcorner)$ is a line bundle. When $M=0$ and $X$ is smooth, the above theorem follows from the traditional Kawamata’s semi-positivity theorem in [K98]. Compared with [K98], we add the extra assumption (1). The condition is required by analysis of the singularities of the metric on $M_{Y}$. However, applying the weak semi-stable reduction theorem in [AK], we can always assume (1) with the birational modification. Then, we need to generalize the theorem for quasi-smooth and toroidal. Therefore, we state the above situation. The proof follows from the idea of the second author’s paper [Ta]. However, by the contribution of $M$, the estimation becomes more complicated than [Ta].
Next, using twisted Kawamata’s semi-positivity theorem 1.1, we prove the following theorem:
Theorem 1.2 ( =Theorem 5.4).
Let $(X,B+M)$ be a projective generalized klt pair such that a $\mathbb{Q}$-line bundle $M^{\prime}$ on a smooth higher birational model $X^{\prime}\to X$ has a semi-positive singular Hermitian metric $h_{M^{\prime}}$ with a vanishing Lelong number for the potential at every point of $X^{\prime}$. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
We collect the notion of generalized pairs by Birkar in Section 5. When $K_{X}+B+M$ is big, the above theorem directly follows from [BCHM] since we perturb $B+M$ with $K_{X}+B+M$ and produce a usual klt boundary $B^{\prime}$ such that $B^{\prime}\sim_{\mathbb{Q}}B+M+\varepsilon(K_{X}+B+M)$ for a small $\varepsilon>0$. Thus, we apply the Fujino–Mori argument [FM]. Finally, we note the interesting difference between the usual and generalized pair in the Kawamata semi-positivity, the canonical bundle formula, and the finite generation. For generalized pairs, we do now know the functorial behavior of the moduli part of [Am04, Theorem 0.2 (1)]. For the case of (generalized) log Calabi–Yau fibrations, it is known by [Fil]. However, in general, it is unknown. Fortunately, this does not matter for our application. Moreover, one may ask the case of log canonical in Theorem 1.2. However, the finite generation of lc generalized canonical rings does not hold by [G, Remark 3.5]. After putting this paper on the arXiv, we know that Hacon and Păun are working with the similar idea with us in [HP]. They also treat various applications for the adjunction theory.
Aknowledgement
The first author was partially supported by grants JSPS KAKENHI ##\##18H01108, 19KK0345, 20H00111, 21H00970, 22H01112, 23H01064. The second author was partially supported by grants JSPS KAENHI ##\##21H00979. The first author thanks Professor Paolo Cascini and Chi-Kang Chang for the discussion.
2. Preliminary
2.1. Algebraic preliminary
We organize some notations and preliminary discussions for our main argument. Let everything be as in 1.1.
Notation 2.1.1.
(1) We write, in general, as $D=D_{+}-D_{-}$ with the positive part $D_{+}=\sum_{d_{i}>0}d_{i}P_{i}$ and the negative part $D_{-}=\sum_{d_{i}<0}(-d_{i})P_{i}$ (without common components). We have $D=D^{h}_{+}-D^{h}_{-}+D^{v}_{+}-D^{v}_{-}$.
(2) We let ${\Delta}_{D}=D_{+}+\ulcorner{D_{-}}\urcorner-D_{-}$. It is effective, and the pair $(X,{\Delta}_{D})$ is klt by 1.1 (2). We see $D={\Delta}_{D}-\ulcorner{D_{-}}\urcorner$, $\ulcorner{-D}\urcorner=\ulcorner{D_{-}}\urcorner$, and hence $K_{X}+D=K_{X}+{\Delta}_{D}-\ulcorner{D_{-}}\urcorner$, $K_{X/Y}+{\Delta}_{D}+M\sim_{{\mathbb{Q}}}\ulcorner{D_{-}}\urcorner+f^{*}L$.
If $D=\sum d_{j}P_{j}$ with $d_{j}\in{\mathbb{Q}}$ in general, $(X,D)$ is sub-klt if and only if all $d_{j}<1$ ($V$-normal crossing support with coefficients $<1$), [KM, 5.20 (4)].
We recall the notion of singularities of pairs.
Definition 2.1.2 (Singularities of pairs).
Let $X$ be a normal variety and let $\Delta$ be an $\mathbb{R}$-divisor on $X$ such that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. Let $\varphi:Y\rightarrow X$ be a log resolution of $(X,\Delta)$. We set
$K_{Y}=\varphi^{*}(K_{X}+\Delta)+\sum a_{i}E_{i},$
where $E_{i}$ is a prime divisor on $Y$ for every $i$. The pair $(X,\Delta)$ is called
  • (a)
    sub Kawamata log terminal ((((subklt, for short)))) if $a_{i}>-1$ for all $i$, or
  • (b)
    sub log canonical ((((sublc, for short)))) if $a_{i}\geq-1$ for all $i$.
If $\Delta$ is effective and $(X,\Delta)$ is subklt (resp. sublc), then we simply call it klt (resp. lc).
Let us recall, in general, a resolution of singularities theorem of morphisms due to Abramovich and Karu [AK]. (We refer [K15, Theorem 2]. See [K15, §1] for terminology.)
Theorem 2.1.3.
Let $f:X\to Y$ be a surjective morphism of smooth projective varieties with connected fibers and $Z$ a closed subset of $X$. Then there exists a quasi-smooth projective toroidal variety $(X^{\prime},B^{\prime})$, a smooth projective toroidal variety $(Y^{\prime},C^{\prime})$, a projective morphism $f^{\prime}:X^{\prime}\to Y^{\prime}$ with connected fibers and projective birational morphisms $v_{X}:X^{\prime}\to X,v_{Y}:Y^{\prime}\to Y$ such that $v_{Y}\circ f^{\prime}=f\circ v_{X}$ and which satisfy the following conditions.
(1) $f^{\prime}:(X^{\prime},B^{\prime})\to(Y^{\prime},C^{\prime})$ is a toroidal morphism.
(2) All the fibers of $f^{\prime}$ have the same dimension ($f^{\prime}$ is flat in fact).
(3) $v_{X}^{-1}(Z)\subset B^{\prime}$.
We call the morphism $f^{\prime}:(X^{\prime},B^{\prime})\to(Y^{\prime},C^{\prime})$ a well-prepared birational model of $f:X\to Y$. We will not use the so-called weak semi-stable reduction. A toroidal variety $(X,B)$ is said to be smooth if $X$ is smooth and $B$ has only normal crossings. It is quasi-smooth if there exists a local toric model of each point, which is a quotient of a smooth toric local model by a finite abelian group (some people say that $X$ is a $V$-manifold and $B$ is $V$-normal crossing).
We will pass to a smooth model of $X$. We may lose the well-preparedness.
Lemma 2.1.4.
Let $f:(X,P)\to(Y,Q),D,M,L$ be as in 1.1 (1)–(4). Let ${\alpha}:X^{\prime}\to X$ be a log-resolution of the pair $(X,P)$, and put $f^{\prime}=f\circ{\alpha}:X^{\prime}\to Y$. Write as
$K_{X^{\prime}}+D^{\prime}={\alpha}^{*}(K_{X}+D)$
for a ${\mathbb{Q}}$-divisor $D^{\prime}$ on $X^{\prime}$, and let $D^{\prime}=\sum_{j}d^{\prime}_{j}P^{\prime}_{j}\ (d^{\prime}_{j}\in{\mathbb{Q}})$ be the decomposition into irreducible components. Let $P^{\prime}=\sum_{d^{\prime}_{j}>0}P^{\prime}_{j}$. Then, $f^{\prime}:(X^{\prime},P^{\prime})\to(Y,Q),D^{\prime},M^{\prime}={\alpha}^{*}M,L$ satisfy the following (1)–(4):
  • (1)
    $f^{\prime}:\mbox{{\rm Supp}}\,{D^{\prime}}^{h}\to Y$ is relatively normal crossing over $Y\setminus Q$ and $f(\mbox{{\rm Supp}}\,{D^{\prime}}^{v})\subseteq Q$, where ${D^{\prime}}^{h}$ and ${D^{\prime}}^{v}$ are the horizontal and vertical part of $D^{\prime}$ respectively,
  • (2)
    $d_{j}<1$,
  • (3)
    $\dim_{\mathbb{C}(\eta)}f^{\prime}_{*}\mathcal{O}_{X^{\prime}}(\ulcorner-D^{\prime}\urcorner)\otimes_{\mathcal{O}_{Y}}\mathbb{C}(\eta)=1$, where $\eta$ is the generic point of $Y$, and
  • (4)
    $K_{X^{\prime}}+D^{\prime}+M^{\prime}\sim_{\mathbb{Q}}{f^{\prime}}^{*}(K_{Y}+L)$.
Moreover, consider the discriminant part ${\Delta}_{Y}^{\prime}$ and the moduli part $M_{Y}^{\prime}=L-{\Delta}_{Y}^{\prime}$ for $f^{\prime}$ as in 1.1 (5). Then, ${\Delta}_{Y}^{\prime}={\Delta}_{Y}$ holds and hence $M_{Y}^{\prime}=M_{Y}$ as a consequence.
Proof.
(i) We can see that $(X^{\prime},D^{\prime})$ is sub-klt as $(X,D)$ is sub-klt, and satisfies (1) and (2). We have $D^{\prime}={\alpha}^{*}D-K_{X^{\prime}/X}=\widetilde{D}_{+}-\widetilde{D}_{-}+E_{+}-E_{-}$, where $\widetilde{D}_{+}$ and $\widetilde{D}_{-}$ are the strict transforms of $D_{+}$ and $D_{-}$ respectively, and where $E_{+}$ and $E_{-}$ are effective ${\alpha}$-exceptional divisors without common components. All the coefficients of $\widetilde{D}_{+}$ and $E_{+}$ need to be strictly less than 1. Then $\ulcorner{-D^{\prime}}\urcorner=\ulcorner{\widetilde{D}_{-}}\urcorner+\ulcorner{E_{-}}\urcorner$. At a general point $y\in Y$, $X_{y}$ is normal, $X^{\prime}_{y}$ is smooth, and the effective divisor $\ulcorner{E_{-}}\urcorner|_{X^{\prime}_{y}}$ is ${\alpha}_{y}$-exceptional (or $\ulcorner{E_{-}}\urcorner|_{X^{\prime}_{y}}=0$) with respect to the induced morphism ${\alpha}_{y}:X^{\prime}_{y}\to X_{y}$. Using this, we have $11\leq h^{0}(X^{\prime}_{y},\ulcorner{\widetilde{D}_{-}}\urcorner+\ulcorner{E_{-}}\urcorner)\leq h^{0}(X_{y},\ulcorner{D_{-}}\urcorner)=$ 1. Thus $f^{\prime}_{*}{\mathcal{O}}_{X^{\prime}}(\ulcorner{-D^{\prime}}\urcorner)$ is of rank 1.
(ii) Let us see that the discriminant part does not depend on the birational model of $X$. For $f:X\to Y$, we set
$b_{Q_{l}}=\max\left\{t\in\mathbb{Q}\,\left|\,\begin{array}[]{l}{\text{$(X,D+tf^{*}{Q_{l}})$ is sublc over}}\\ {\text{the generic point of $Q_{l}$}}.\end{array}\right.\right\}.$
Let us see first that $\delta_{l}=1-b_{Q_{l}}$. By shrinking $Y$ at the generic point of $Q_{l}$, we may assume that all $P_{j}$ dominate $Q_{l}$. We consider the pair $(X,D+(1-\delta_{l})f^{*}Q_{l})$. Then, the pair is quasi-smooth, and every coefficient is less than or equal to $1$. Moreover, there exists some $P_{j}$ such that the coefficient is $1$. Then, it is enough to show that for a quasi-smooth pair $(X,B)$ is sublc if and only if $B\leq 1$ (i.e., every coefficient of $B$ is less than or equal to 1). But this follows from [KM, Proposition 5.20 (4)].
We let
${f^{\prime}}^{*}Q_{l}$ $=$ $\sum_{j}w^{\prime}_{lj}P^{\prime}_{j},\ \text{where}\ w^{\prime}_{lj}>0,$
$\bar{d^{\prime}}_{j}$ $=$ $\frac{d^{\prime}_{j}+w^{\prime}_{lj}-1}{w^{\prime}_{lj}}\ \text{if}\ f^{\prime}(P^{\prime}_{j})=Q_{l},$
$\delta^{\prime}_{l}$ $=$ $\mathrm{max}\{\bar{d^{\prime}_{j}}|f^{\prime}(P^{\prime}_{j})=Q_{l}\}.$
We would like to show that $\delta_{l}=\delta^{\prime}_{l}$. We observed that $\delta_{l}=1-b_{Q_{l}}$. On the other hand, it holds that $\delta^{\prime}_{l}=1-\mathrm{lct}_{{f^{\prime}}^{*}Q_{lgen}}(X^{\prime},D^{\prime})$, where
$\mathrm{lct}_{{f^{\prime}}^{*}Q_{lgen}}(X^{\prime},D^{\prime})=\max\left\{t\in\mathbb{Q}\,\left|\,\begin{array}[]{l}{\text{$(X^{\prime},D^{\prime}+t{f^{\prime}}^{*}{Q_{l}})$ is sublc over}}\\ {\text{the generic point of $Q_{l}$}}.\end{array}\right.\right\}.$
Since such log-canonical thresholds do not change by the crepant pullback, we have $b_{Q_{l}}=\mathrm{lct}_{{f^{\prime}}^{*}Q_{lgen}}(X^{\prime},D^{\prime})$. That is the desired conclusion. ∎
2.2. Analytic preliminary
We recall a few basic facts on the Lelong number, which measures the “analytic” singularities of plurisubharmonic (psh for short) functions. We refer to [D12, 2.B, 5.B] for some properties of psh functions related to algebraic geometry. We also refer to [D12] for a general theory of singular Hermitian metrics, multiplier ideal sheaves, and so on.
Definition 2.2.1.
Let $\varphi$ be a (quasi-)psh function on an open set $B$ in $\mathbb{C}^{n}$. The Lelong number $\nu(\varphi,x)$ of $\varphi$ at $x\in B$ is defined by
$\nu(\varphi,x)=\liminf_{z\to x}\frac{\varphi(z)}{\mathrm{log}|z-x|}.$
The Lelong number does not depend on the choice of local coordinates and can, hence, be defined on complex manifolds.
A singular Hermitian metric $h$ on a line bundle $L$ can be written as $h=e^{-{\varphi}}$ with ${\varphi}\in L^{1}_{loc}(U,{\mathbb{R}})$ on every local trivialization $L|_{U}\cong U\times{\mathbb{C}}$. Every such ${\varphi}$ is said to be a weight function of $h$. If the curvature current of $h$ is semi-positive, then ${\varphi}$ is psh on $U$. The curvature positivity of $h$ relates to positivity notions in algebraic geometry. The following result is particularly related to this paper, which is a consequence of the regularization theory of singular metrics of Demailly, [D92, Proposition 3.7].
Proposition 2.2.2.
A line bundle on a smooth projective variety admitting a singular Hermitian metric $h$ with semi-positive curvature and whose local weight functions have vanishing Lelong numbers everywhere is nef. Namely, the intersection number with any curve is semi-positive.
The following is a particular case of [Fa, Theorem 2].
Lemma 2.2.3.
Let $f:U\to V$ be a surjective holomorphic map of complex manifolds, and let ${\varphi}$ be a psh function on $V$. Let $p\in U$. Then, $\nu({\varphi},f(p))=0$ if and only if $\nu(f^{*}{\varphi},p)=0$.
Although we do not try to define the Lelong number on singular spaces, we can introduce it.
Definition 2.2.4.
Let $V$ be a normal complex space, and let ${\varphi}$ be a psh function on $V$. Let $q\in V$. The function ${\varphi}$ has zero (or vanishing) Lelong number at $q$, if there exist a complex manifold $U$, a surjective holomorphic map $f:U\to V$ and a point $p\in f^{-1}(q)$ such that $\nu(f^{*}{\varphi},p)=0$.
This property is preserved by an arbitrary pull-back.
Lemma 2.2.5.
Let $V$ be a normal complex space, and let ${\varphi}$ be a psh function on $V$. Let $q\in V$. Suppose ${\varphi}$ has zero Lelong number at $q$. Let $g:W\to V$ be a surjective holomorphic map from a normal complex space $W$ and take a point $r\in g^{-1}(q)$. Then, $g^{*}{\varphi}$ has zero Lelong number at $r$. In particular, if $W$ is smooth, $\nu(g^{*}{\varphi},r)=0$.
Proof.
Let $f:U\to V$ and $p\in f^{-1}(q)$ be as in 2.2.4 such that $\nu(f^{*}{\varphi},p)=0$. (We can restrict everything over a relatively compact open neighborhood of $q$.) By taking the fiber product $W\times_{V}U$ and a resolution of singularities of it, we can find a complex manifold $X$ with surjective holomorphic maps $f^{\prime}:X\to W$ and $g^{\prime}:X\to U$ such that $g\circ f^{\prime}=f\circ g^{\prime}$. We can also take a point $x\in X$ such that $f^{\prime}(x)=r$ and $g^{\prime}(x)=p$. By 2.2.3, the property $\nu(f^{*}{\varphi},p)=0$ is equivalent to $\nu((g\circ f^{\prime})^{*}{\varphi},x)=0$, which is saying that $g^{*}{\varphi}$ has zero Lelong number at $r$. ∎
We will use the following criterion to check the vanishing of a Lelong number.
Lemma 2.2.6.
Let $U$ be a unit ball in $({\mathbb{C}}^{n},z=(z_{1},\ldots,z_{n}))$, and let ${\varphi}$ and $u$ be quasi-psh functions on $U$ with $\nu(u,0)=n$. Let $k\in(0,2]$. Suppose $e^{-{\varphi}}e^{-cu}\in L^{1}_{loc}(U)$ for any $c\in(0,k)$. Then $\nu({\varphi},0)\leq(2-k)n$. For example, we can conclude $\nu({\varphi},0)=0$, if we can take $k=2$ and $u=\log|\prod_{j=1}^{n}z_{j}|$.
Proof.
Suppose on the contrary that $\nu({\varphi},0)>(2-k)n$. We let $a>0$ so that $\nu({\varphi},0)=(2-k)n+a$, and take $b>0$ such that $b<\min\{k,a/n\}$. Then ${\varphi}+(k-b)u$ is quasi-psh on $U$ and $\nu({\varphi}+(k-b)u,0)\geq\nu({\varphi},0)+(k-b)\nu(u,0)=2n+a-bn>2n$. Then $e^{-{\varphi}}e^{-(k-b)u}\not\in L^{1}_{loc}(U)$ ([D12, 5.6]). This is a contradiction. ∎
Psh functions with vanishing Lelong numbers do not affect multiplier ideals.
Lemma 2.2.7.
Let $U$ be a unit ball in ${\mathbb{C}}^{n}$, and let ${\varphi}$ and $u$ be quasi-psh functions on $U$. Suppose $\nu({\varphi},0)=0$. Then ${\mathcal{I}}(e^{-{\varphi}}e^{-u})={\mathcal{I}}(e^{-u})$ on a neighborhood of 00.
Here, ${\mathcal{I}}(e^{-u})\subset{\mathcal{O}}_{U}$ is a multiplier ideal sheaf. It is the sheaf of germs of holomorphic functions $f$ which is $L^{2}$ with respect to $e^{-u}dv$, where $dv$ is the standard Lebesgue measure on $U$, namely $|f|^{2}e^{-u}\in L^{1}_{loc}(U)$. The multiplier ideal sheaf is coherent if $u$ is quasi-psh ([D12, 5.7]).
Proof.
${\mathcal{I}}(e^{-{\varphi}}e^{-u})\subset{\mathcal{I}}(e^{-u})$ is clear. We take $f\in{\mathcal{I}}(e^{-u})_{0}$. We can take a relatively compact open neighborhood $V$ of 00 in $U$ such that $\int_{V}|f|^{2}e^{-u}<\infty$. Here, we take the Lebesgue measure as a volume form on $U$. We take $p,q\in{\mathbb{R}}_{>1}$ with $11/p+1/q=$ 1 such that ${\mathcal{I}}(e^{-qu})={\mathcal{I}}(e^{-u})$ on $V$. The last property is the so-called strong openness property of multiplier ideals [GZ] (see also [Be]). As $\nu(p{\varphi},0)=0$, we can take a neighborhood $W\subset V$ of 00 such that $e^{-p{\varphi}}\in L^{1}(W)$ by Skoda’s lemma ([D12, 5.6]). Then, by Hölder’s inequality,
$\int_{W}|f|^{2}e^{-{\varphi}}e^{-u}\leq\Big{(}\int_{W}|f|^{2}e^{-p{\varphi}}\Big{)}^{1/p}\Big{(}\int_{W}|f|^{2}e^{-q{\varphi}}\Big{)}^{1/q}.$
Thus, we have $f\in{\mathcal{I}}(e^{-{\varphi}}e^{-u})_{0}$. If $u$ has only analytic singularities, the openness property is an elementary property. ∎
The next lemma is trivial.
Lemma 2.2.8.
Let $a_{j},b_{j}\in{\mathbb{R}}\ (j=1,2,\ldots,n)$ with $a_{j}<2$ for any $j$. Then
$\prod_{j=1}^{n}\frac{(-\log|z_{j}|)^{b_{j}}}{|z_{j}|^{a_{j}}}\in L^{1}_{loc}({\mathbb{C}}^{n}).$
3. Fiber integral
3.1. Basic estimate: Smooth total space
Let everything be as in 1.1, and suppose $X$ is smooth. We let $\dim X=n+m$ and $\dim Y=m$. We explain a key object: a fiber integral, which will appear repeatedly in this paper (and can be formulated in various other situations).
3.1.1.
We let
$E=\ulcorner{D_{-}}\urcorner-K_{X/Y},$
which is a line bundle on $X$. We take any Hermitian metric $h$ on $E$. We take a (root of) section $s_{{\Delta}_{D}}$ whose divisor is ${\Delta}_{D}$, and take a Hermitian metric $h_{{\Delta}_{D}}$ on the ${\mathbb{Q}}$-line bundle ${\Delta}_{D}$. Let ${\varphi}\in L^{1}_{loc}(X,{\mathbb{R}})$ be a quasi-psh function. We finally let
$g=h\ \frac{e^{-{\varphi}}}{|s_{{\Delta}_{D}}|_{h_{{\Delta}_{D}}}^{2}}$
be a singular Hermitian metric on $E$. We take a section
${\sigma}=s_{\ulcorner{D_{-}}\urcorner}\in H^{0}(X,\ulcorner{D_{-}}\urcorner)=H^{0}(X,K_{X/Y}+E)$
whose divisor is $\ulcorner{D_{-}}\urcorner=\ulcorner{-D}\urcorner$. We regard ${\sigma}$ as an $E$-valued relative holomorphic $n$-form. We then consider the associated relative (possibly singular) volume form $\omega$ and the fiber ($L^{2}$-)integral of ${\sigma}$ with respect to $g$:
$\omega=g\ c_{n}{\sigma}\wedge\overline{\sigma},\ {\rm and}\ F_{\omega}(y)=\int_{X_{y}}\omega\ \in[0,+\infty]$
for $y\in Y\setminus Q$. Here $c_{n}=\sqrt{-1}^{n^{2}}$ so that the pairing becomes semi-positive on each fiber. ∎
For $y\in Y\setminus Q$, $F_{\omega}(y)<+\infty$ if and only if ${\sigma}|_{X_{y}}\in H^{0}(X_{y},K_{X_{y}}\otimes E|_{X_{y}}\otimes{\mathcal{I}}(g|_{X_{y}}))$, where ${\mathcal{I}}(g|_{X_{y}})$ is the multiplier ideal sheaf of $g|_{X_{y}}$. We understand $F_{\omega}(y)=+\infty$ if ${\varphi}|_{X_{y}}\equiv-\infty$ on a fiber for example. We know, in general, $F_{\omega}(y)$ is lower-semi-continuous on which $f$ is smooth, [HPS, 22.5].
We are interested in the behavior of $F_{\omega}(y)$ as $y$ approaches to $Q$. Here, we are only concerned about the singularity type of $g$, not about the curvature property of $g$, e.g.  semi-positivity. The choices of smooth metrics $h,h_{{\Delta}_{D}}$ also do not affect the following conclusion.
Theorem 3.1.2.
Suppose $X$ is smooth and ${\Delta}_{0}=0$ in 1.1, and use the notations in 3.1.1. Suppose the Lelong number $\nu({\varphi},x)=0$ for every $x\in X$. Let $y_{0}\in Y$ and let $V\subset({\mathbb{C}}^{m},t=(t_{1},\ldots,t_{m}))$ be a local coordinate centered at a point $y_{0}\in Y$. Then
(1) Upper estimate. $F_{\omega}/|\prod_{j=1}^{m}t_{j}|^{c}\in L^{1}_{loc}(V,0)$ for any $c\in(0,2)$. In particular, if $-\log F_{\omega}$ is quasi-psh on a neighborhood of $y_{0}$, one has $\nu(-\log F_{\omega},y_{0})=0$.
(2) Lower estimate. There exist positive constants ${\delta}(<1)$ and $c_{1}$ such that, for any $t\in V\setminus Q$ with $|t|<{\delta}$, $F_{\omega}(t)\geq c_{1}$ holds.
Proof.
(1) Step 1. Preliminary discussion. We take $c\in(0,2)$. We put $T=\prod_{j=}^{m}t_{j}$, and $|s_{{\Delta}_{D}}|=|s_{{\Delta}_{D}}|_{h_{{\Delta}_{D}}}$ for short. We note $hc_{n}{\sigma}\wedge\overline{\sigma}|dt|^{2}$ is a $C^{\infty}$ semi-positive volume form on $X_{V}=f^{-1}(V)$, where $dt=dt_{1}\wedge\cdots\wedge dt_{m}$ and $|dt|^{2}=c_{m}dt\wedge d\overline{t}$. Then
$\int_{V}F_{\omega}\ \frac{|dt|^{2}}{|T|^{c}}=\int_{X_{V}}\frac{e^{-{\varphi}}}{|s_{{\Delta}_{D}}|^{2}|T|^{c}}hc_{n}{\sigma}\wedge\overline{\sigma}|dt|^{2}.$
We take an open covering ${\mathcal{U}}=\{U_{{\lambda}}\}_{{\lambda}}$ of $X$, which is fine enough and locally finite, and take a partition of unity $\{\rho_{{\lambda}}\}_{{\lambda}}$ subordinate to ${\mathcal{U}}$. We can suppose $\overline{X_{V}}\cap U_{{\lambda}}\neq\emptyset$ for a finite number of ${\lambda}$. Then
$\int_{V}F_{\omega}\ \frac{|dt|^{2}}{|T|^{c}}=\sum_{{\lambda}}\int_{U_{{\lambda}}\cap X_{V}}\rho_{{\lambda}}\frac{e^{-{\varphi}}}{|s_{{\Delta}_{D}}|^{2}|T|^{c}}hc_{n}{\sigma}\wedge\overline{\sigma}|dt|^{2}.$
Thus, it is enough to evaluate an integral on each local coordinate $U$ of $X$. We let $X_{t}=f^{-1}(t)$ for $t\in V$.
Step 2. Local coordinate. At each $x_{0}\in X_{0}$, we can find local coordinates $U\subset({\mathbb{C}}^{n+m},x=(x_{1},\ldots,x_{n+m})),V\subset({\mathbb{C}}^{m},t=(t_{1},\ldots,t_{m}))$ (we suppose $|x_{i}|\leq 1$ for all $i$, $|t_{j}|\leq 1$ for all $j$), such that $f|_{U}:U\to V$ is given by $t=f(x),t_{j}=f_{j}(x),j=1,\ldots,m$, with
$t_{1}=\prod_{j=1}^{\ell_{1}}x_{j}^{a_{j}},\ t_{2}=\prod_{j=\ell_{1}+1}^{\ell_{2}}x_{j}^{a_{j}},\ \ldots,\ t_{m}=\prod_{j=\ell_{m-1}+1}^{\ell_{m}}x_{j}^{a_{j}}$
for some $0=\ell_{0}<\ell_{1}<\ldots<\ell_{m}\leq n+m$ and $a_{j}\in{\mathbb{Z}}_{>0}$. (The coordinates $x,t$ depend on $x_{0}$, but) Moreover $Q\cap V\subset\{t_{1}\cdots t_{m}=0\}$ and $P\cap U\subset\{x_{1}\cdots x_{n+m}=0\}$. This $V\subset({\mathbb{C}}^{m},t)$ may be different from the one we took in the statement of 3.1.2. However, it does not matter what our purpose is. We can write as $\ell_{i}=\ell_{i-1}+n_{i}+1$ with $n_{i}\in{\mathbb{Z}}_{\geq 0}$ and $\sum_{i=1}^{m+1}n_{i}=n$, where $n_{m+1}:=n+m-\ell_{m}$. We denote by $y_{j}=x_{\ell_{j}},b_{j}=a_{\ell_{j}}$ for $j=1,\ldots,m$, and by
$\mathbf{x}_{j}=(x_{\ell_{j-1}+1},\ldots,x_{\ell_{j-1}+n_{j}}),\ \mathbf{a}_{j}=(a_{\ell_{j-1}+1},\ldots,a_{\ell_{j-1}+n_{j}}),$
$\mathbf{x}_{m+1}=(x_{\ell_{m}+1},\ldots,x_{n+m}).$
We will use a multi-index convention. For example $t_{j}=\mathbf{x}_{j}^{\mathbf{a}_{j}}y_{j}^{b_{j}}$.
Step 3. Integrand. We may suppose that every (${\mathbb{Q}}$-)line bundle appearing in our situation is trivialized on a neighborhood of $\overline{U}$ or of $\overline{V}$. We take a local frame $dx=\wedge_{i=1}^{n+m}dx_{i}$ of $K_{X}|_{U}$ and $dt=\wedge_{j=1}^{m}dt_{j}$ (as before) of $K_{Y}|_{V}$ respectively. Using relations $\frac{dt_{j}}{t_{j}}=\sum_{i=\ell_{j-1}+1}^{\ell_{j-1}+n_{j}}a_{i}\frac{dx_{i}}{x_{i}}+b_{j}\frac{dy_{j}}{y_{j}}$, we have $\wedge_{i=\ell_{j-1}+1}^{\ell_{j}}dx_{i}=\frac{y_{j}}{b_{j}t_{j}}\big{(}\wedge_{i=\ell_{j-1}+1}^{\ell_{j-1}+n_{j}}dx_{i}\big{)}\wedge dt_{j}$, in other term $d\mathbf{x}_{j}\wedge dy_{j}=\frac{y_{j}}{b_{j}t_{j}}d\mathbf{x}_{j}\wedge dt_{j}$. Then, a local frame $\frac{dx}{dt}$ of $K_{X/Y}$ on $U$ can be expressed as
$\frac{dx}{dt}=\bigwedge_{j=1}^{m}\Big{(}\frac{y_{j}}{b_{j}t_{j}}d\mathbf{x}_{j}\Big{)}\wedge d\mathbf{x}_{m+1},$
which satisfies $\frac{dx}{dt}\wedge f^{*}dt=\pm dx$. We also let $|dx|^{2}=c_{n+m}dx\wedge d\overline{x}$, and $\big{|}\frac{dx}{dt}\big{|}=c_{n}\frac{dx}{dt}\wedge\overline{\frac{dx}{dt}}$.
Regarding ${\sigma}\in H^{0}(X,K_{X/Y}+E)={\rm Hom}_{{\mathcal{O}}_{X}}(f^{*}K_{Y},K_{X}+E)$, we have ${\sigma}(f^{*}dt)|_{U}={\sigma}_{U}dx=\pm{\sigma}_{U}\frac{dx}{dt}\wedge f^{*}dt$ for some ${\sigma}_{U}\in H^{0}(U,{\mathcal{O}}_{X})(\cong H^{0}(U,E))$, and hence
${\sigma}|_{X_{t}}=\pm{\sigma}_{U}\frac{dx}{dt}\Big{|}_{X_{t}}$
for $t\in V\setminus Q$. The metric $h$ can be seem as a function in $C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$. Thus we have
$hc_{n}{\sigma}\wedge\overline{\sigma}|dt|^{2}=A_{U}|{\sigma}_{U}|^{2}|dx|^{2}$
on $U$ for some $A_{U}\in C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$. We finally take a cut-off function $\rho_{U}\in C^{\infty}_{c}(U,{\mathbb{R}}_{\geq 0})$. Then, in view of Step 1, we need to estimate
$I_{U}=\int_{U}\rho_{U}e^{-{\varphi}}\frac{A_{U}|{\sigma}_{U}|^{2}}{|T|^{c}|s_{{\Delta}_{D}}|^{2}}|dx|^{2}.$
Step 4. Reduction to an un-twisted case. We introduce auxiliary quantities. Recall $f^{*}Q_{j}=\sum_{i}a_{ji}P_{i}$ for each $j$. We set $c_{j}=\max_{i}\{d_{i}/a_{ji}\ {\rm if}\ f(P_{i})=Q_{j}\}$, and $\widetilde{c}=\max_{j}c_{j}\in{\mathbb{R}}$. We take $\varepsilon>0$ such that $c(1+\varepsilon)+2\widetilde{c}\varepsilon<2$. A simpler condition $c(1+\varepsilon)+2d_{i}\varepsilon<2$ for any coefficient $d_{i}$ of $P_{i}$ also works (which in fact implies $c(1+\varepsilon)+2\widetilde{c}\varepsilon<2$). We also require that $d_{i}(1+\varepsilon)<1$ holds for any $i$ with $d_{i}<1$. We then take $q=1+\varepsilon$ and $p\gg 1$ such that $\frac{1}{p}+\frac{1}{q}=1$.
By Hölder’s inequality, we have
$I_{U}\leq\big{\|}\rho_{U}^{1/p}e^{-{\varphi}}\big{\|}_{L^{p}(U,|dx|^{2})}\cdot\Big{\|}\rho_{U}^{1/q}\frac{A_{U}|{\sigma}_{U}|^{2}}{|T|^{c}|s_{{\Delta}_{D}}|^{2}}\Big{\|}_{L^{q}(U,|dx|^{2})}.$
The first term on the right hand side is finite:
$(J_{p}:=)\ \big{\|}\rho_{U}^{1/p}e^{-{\varphi}}\big{\|}_{L^{p}(U,|dx|^{2})}^{p}=\int_{U}\rho_{U}e^{-p{\varphi}}|dx|^{2}<\infty,$
by virtue of our assumption $\nu({\varphi},x)=0$ for any $x\in X$. We next consider
$J_{q}:=\int_{U}\rho_{U}\frac{A_{U}^{q}|{\sigma}_{U}|^{2q}}{|T|^{cq}|s_{{\Delta}_{D}}|^{2q}}|dx|^{2}.$
By Fubini, we have
$J_{q}=\int_{t\in V}\frac{|dt|^{2}}{|T|^{cq}}\int_{X_{t}\cap U}\rho_{U}A_{U}^{q}\frac{|{\sigma}_{U}|^{2q}}{|s_{{\Delta}_{D}}|^{2q}}\Big{|}\frac{dx}{dt}\Big{|}^{2}.$
We will show in the next step that the fiber integral is estimated as
$\int_{X_{t}\cap U}\rho_{U}A_{U}^{q}\frac{|{\sigma}_{U}|^{2q}}{|s_{{\Delta}_{D}}|^{2q}}\Big{|}\frac{dx}{dt}\Big{|}^{2}\leq\frac{A}{|T|^{2\widetilde{c}\varepsilon}}\prod_{j=1}^{m}(-\log|t_{j}|)^{n}$
for any $t\in V\setminus Q$, where $A>0$ is a constant independent of $t$. This estimate deduces that
$J_{q}\leq A\int_{t\in V}\frac{\prod_{j=1}^{m}(-\log|t_{j}|)^{n}}{|T|^{cq+2\widetilde{c}\varepsilon}}\ |dt|^{2}<\infty.$
The last integrability follows from $cq+2\widetilde{c}\varepsilon<2$ and 2.2.8. We have $I_{U}<\infty$ as a conclusion.
The statement $\nu(-\log F_{\omega},y)=0$ is a consequence of this integrability combined with 2.2.6.
Step 5. Estimate of fiber integral. We can suppose $V\cap Q=\{t_{1}\cdots t_{k}=0\}$ for some $k$ with $0\leq k\leq m$, and $f$ is smooth over $V\setminus Q$ and singular over $V\cap Q$ (cf. [Ta, 2.2 (2)]). We can let $Q_{j}\cap V=\{t_{j}=0\}$ for $1\leq j\leq k$, and $P_{i}\cap U=\{x_{i}=0\}$ for $1\leq i\leq\ell_{k}$ (such that $P_{i}$ is $f$-vertical and appeared in $D$). For a notational simplicity, we suppose $k=m$ (and hence $\ell_{k}=\ell_{m}$). Other cases $10\leq k\leq m-$ 1 can be deduced from the case $k=m$. In particular $P_{i}$ is $f$-vertical if and only if $1\leq i\leq\ell_{m}$. We write as
$D^{h}={\rm div}\Big{(}\prod_{i>\ell_{m},d_{i}<1}x_{i}^{d_{i}}\Big{)},\ D^{v}_{-}={\rm div}\Big{(}\prod_{i\leq\ell_{m},-d_{i}>0}x_{i}^{-d_{i}}\Big{)},\ D^{v}_{+}={\rm div}\,(1)=0.$
Here $D_{+}^{v}=0$ by our assumption: $f$ is equi-dimensional and ${\Delta}_{Y}=0$ (cf. [Ta, 2.2 (3)]).
We estimate the integrand of the fiber integral in $J_{q}$. We have
$|{\sigma}_{U}|^{2}$ $=A_{1}\cdot\prod_{i,-d_{i}>0}|x_{i}|^{2\ulcorner{-d_{i}}\urcorner},$
$|s_{{\Delta}_{D}}|^{2}$ $=A_{2}\cdot\prod_{i>\ell_{m},0<d_{i}<1}|x_{i}|^{2d_{i}}\prod_{i,-d_{i}>0}|x_{i}|^{2(\ulcorner{-d_{i}}\urcorner+d_{i})}$
on $U$, where $A_{1},A_{2}\in C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$. Then
$\frac{|{\sigma}_{U}|^{2q}}{|s_{{\Delta}_{D}}|^{2q}}\Big{|}\frac{dx}{dt}\Big{|}^{2}=A_{3}\bigwedge_{j=1}^{m+1}W_{j}$
holds with $A_{3}=A_{1}^{q}A_{2}^{q}\prod_{j=1}^{m}b_{j}^{-2}\in C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$, and
$W_{j}=(|y_{j}|^{2})^{-d_{\ell_{j}}q+1-a_{\ell_{j}}}\prod_{i=\ell_{j-1}+1}^{\ell_{j-1}+n_{j}}(|x_{i}|^{2})^{-d_{i}q-a_{i}}\ c_{n_{j}}d\mathbf{x}_{j}\wedge d\overline{\mathbf{x}}_{j}\ \ (j=1,\ldots,m),$
$W_{m+1}=\frac{\prod_{i>\ell_{m},d_{i}<0}|x_{i}|^{-2d_{i}q}}{\prod_{i>\ell_{m},0<d_{i}<1}|x_{i}|^{2d_{i}q}}\ c_{n_{m+1}}d\mathbf{x}_{m+1}\wedge d\overline{\mathbf{x}}_{m+1}.$
We now start to estimate. Let us look at $W_{1}$ for example, and show
$W_{1}\leq\frac{1}{|t_{1}|^{2c_{1}\varepsilon}}\frac{c_{n_{1}}d\mathbf{x}_{1}\wedge d\overline{\mathbf{x}}_{1}}{|\mathbf{x}_{1}|^{2}}$
as $(n_{1},n_{1})$-form on $U$, where $|\mathbf{x}_{1}|^{2}=\prod_{i=1}^{n_{1}}|x_{i}|^{2}$ (and $|\mathbf{x}_{j}|^{2}=\prod_{i=\ell_{j-1}+1}^{\ell_{j-1}+n_{j}}|x_{i}|^{2}$ in general).
(i) For $t_{1}=\prod_{j=1}^{\ell_{1}}x_{j}^{a_{j}}$, we have $|t_{1}|^{c_{1}}=\prod_{j=1}^{\ell_{1}}|x_{j}^{a_{j}}|^{c_{1}}\leq\prod_{j=1}^{\ell_{1}}|x_{j}^{a_{j}}|^{d_{j}/a_{j}}$ (by definition of $c_{1}$ and as $|x_{j}|\leq 1$). Thus, $|t_{1}|^{c_{1}\varepsilon}\leq\prod_{j=1}^{\ell_{1}}|x_{j}|^{d_{j}\varepsilon}$.
(ii) We note, as a consequence of ${\Delta}_{Y}=0$, that for every $j\in\{1,\ldots,m\}$ and $i\in\{\ell_{j-1}+1,\ldots,\ell_{j}\},-d_{i}-a_{i}\geq-1$ holds, in particular $-d_{\ell_{j}}+1-a_{\ell_{j}}\geq 0$ holds. Recalling $q=1+\varepsilon$, we use it as $-d_{\ell_{j}}q+1-a_{\ell_{j}}\geq-d_{\ell_{j}}\varepsilon$ and $-d_{i}q-a_{i}\geq-1-d_{i}\varepsilon$.
(iii) The “coefficient function” of $W_{1}$ can be estimated as $(|y_{1}|^{2})^{-d_{\ell_{1}}q+1-a_{\ell_{1}}}\prod_{i=1}^{n_{1}}(|x_{i}|^{2})^{-d_{i}q-a_{i}}$ $\leq_{{\rm(ii)}}$ $(|y_{1}|^{2})^{-d_{\ell_{1}}\varepsilon}\prod_{i=1}^{n_{1}}|x_{i}|^{-2}(|x_{i}|^{2})^{-d_{i}\varepsilon}\leq_{{\rm(i)}}|t_{1}|^{-2c_{1}\varepsilon}|\mathbf{x}_{1}|^{-2}$. This gives the desired estimate of $W_{1}$.
We have a similar estimate for $W_{j}\ (j=2,\ldots,m)$. We note $\prod_{j=1}^{m}|t_{j}|^{-2c_{j}\varepsilon}\leq|T|^{-2\widetilde{c}\varepsilon}$. For $W_{m+1}$, we simply estimate it as
$W_{m+1}\leq\frac{c_{n_{m+1}}d\mathbf{x}_{m+1}\wedge d\overline{\mathbf{x}}_{m+1}}{\prod_{i>\ell_{m},0<d_{i}<1}|x_{i}|^{2d_{i}q}}.$
We introduce a (possibly singular) semi-positive $(n,n)$-form
$\psi=\bigwedge_{j=1}^{m}\frac{c_{n_{j}}d\mathbf{x}_{j}\wedge d\overline{\mathbf{x}}_{j}}{|\mathbf{x}_{j}|^{2}}\wedge\frac{c_{n_{m+1}}d\mathbf{x}_{m+1}\wedge d\overline{\mathbf{x}}_{m+1}}{\prod_{i>\ell_{m},0<d_{i}<1}|x_{i}|^{2d_{i}q}}$
on $({\mathbb{C}}^{n+m},x)$. We then have a constant $A_{4}>0$ (independent of $t$) such that
$\rho_{U}A_{U}^{q}\frac{|{\sigma}_{U}|^{2q}}{|s_{{\Delta}_{D}}|^{2q}}\Big{|}\frac{dx}{dt}\Big{|}^{2}=\rho_{U}A_{U}^{q}A_{3}\bigwedge_{j=1}^{m+1}W_{j}\leq\frac{A_{4}}{|T|^{2\widetilde{c}\varepsilon}}\ \psi$
holds on $U$.
As $d_{i}q<1$ in the expression of $\psi$, we can directly apply the way of computation in [Ta, p.1733, Step 4] to conclude that
$\int_{X_{t}\cap U}\psi\leq A_{5}\prod_{j=1}^{m}(-\log|t_{j}|)^{n}$
holds for any $t\in V\setminus Q$, where $A_{5}>0$ is a constant independent of $t$. Thus we have
$\int_{X_{t}\cap U}\rho_{U}A_{U}^{q}\frac{|{\sigma}_{U}|^{2q}}{|s_{{\Delta}_{D}}|^{2q}}\Big{|}\frac{dx}{dt}\Big{|}^{2}\leq\frac{A_{4}A_{5}}{|T|^{2\widetilde{c}\varepsilon}}\prod_{j=1}^{m}(-\log|t_{j}|)^{n}$
for any $t\in V\setminus Q$. This is what we wanted.
(2) Lower estimate. As ${\varphi}$ is quasi-psh, the function $e^{-{\varphi}}$ has a positive lower bound on every compact set of $X$. In particular there exists a constant $c_{0}>0$ such that $e^{-{\varphi}}>c_{0}$ on a neighborhood of $X_{y_{0}}$. (We merely use the existence of a positive lower bound of $e^{-{\varphi}}$, and we do not use the Lelong number condition of ${\varphi}$.) The integrand is
$\omega=h\frac{e^{-{\varphi}}}{|s_{{\Delta}_{D}}|^{2}}c_{n}{\sigma}\wedge\overline{\sigma}\geq h\frac{c_{0}}{|s_{{\Delta}_{D}}|^{2}}c_{n}{\sigma}\wedge\overline{\sigma}.$
The last term is “an un-twisted case”. We then just follow the argument in [Ta, p.1735, Step 5]. ∎
Remark 3.1.3.
Our argument in 3.1.2 is local over $Y$. It can be generalized as follows.
Let $X$ and $Y$ be complex manifolds, $f:X\to Y$ be a proper surjective holomorphic map with connected fibers, which satisfies the conditions (1)–(3) in 1.1 in an appropriate sense for (${\mathbb{Q}}$-)divisors $P,Q$ and $D$. (The property (4), $M,M_{Y}$ and $L$ are irrelevant.) Suppose ${\Delta}_{Y}=0$ in 1.1 (5). Take ${\varphi}\in L^{1}_{loc}(X,{\mathbb{R}})$, $s_{{\Delta}_{D}}$, ${\sigma}=s_{\ulcorner{D_{-}}\urcorner}\in H^{0}(X,\ulcorner{D_{-}}\urcorner)$, and $h,h_{{\Delta}_{D}}$, and construct a sigular Hermitian metric $g$ as in 3.1.1. Put
$F_{\omega}(y)=\int_{X_{y}}g\,c_{n}{\sigma}\wedge\overline{\sigma}\ \in[0,+\infty]$
for $y\in Y\setminus Q$. Let $y_{0}\in Y$ and let $V\subset({\mathbb{C}}^{m},t=(t_{1},\ldots,t_{m}))$ be a local coordinate centered at a point $y_{0}\in Y$.
(1) Suppose ${\varphi}$ is quasi-psh on a neighborhood of $X_{y_{0}}$ and the Lelong number $\nu({\varphi},x)=0$ for every $x\in X_{y_{0}}$. Then, $F_{\omega}/|\prod_{j=1}^{m}t_{j}|^{c}\in L^{1}_{loc}(V,0)$ for any $c\in(0,2)$.
(2) Suppose ${\varphi}$ is locally bounded from above on a neighborhood of $X_{y_{0}}$. Then, there exist positive constants ${\delta}(<1)$ and $c_{1}$ such that, for any $t\in V\setminus Q$ with $|t|<{\delta}$, $F_{\omega}(t)\geq c_{1}$ holds. ∎
3.2. Quasi-smooth total space
We explain fiber integrals in the orbi-fold setting. We regard $\omega_{X}={\mathcal{O}}_{X}(K_{X})$ as a reflexive sheaf of rank 1 and an orbi-line bundle, for example.
3.2.1Fiber integral.
Let $X$ be a complex orbi-fold, $Y$ be a complex manifold, and let $f:X\to Y$ be a proper surjective holomorphic map with connected fibers. We let $\dim X=n+m$ and $\dim Y=m$.
(1) Let $E$ be an orbi-line bundle on $X$, and suppose $E$ is a ${\mathbb{Q}}$-line bundle. Suppose there is a section
${\sigma}\in H^{0}(X,K_{X/Y}+E).$
To be more precise, $K_{X/Y}+E$ means the double dual of the tensor product of orbi-line bundles $K_{X/Y}$ and $E$. We regard ${\sigma}$ as an orbi-line bundle $E$-valued relative holomorphic $n$-form on $X$ in the following sense. For every $x\in X$, we can take an open neighborhood $U\subset X$ of $x$ such that there is an open neighborhood $\widetilde{U}$ of $({\mathbb{C}}^{n+m},0)$ and a finite group $G$, which is “small”, acting on $\widetilde{U}$ and $U$ is the quotient $U=\widetilde{U}/G$. Let $\pi:\widetilde{U}\to U=\widetilde{U}/G$ be the quotient map. We have $K_{\widetilde{U}}=\pi^{*}K_{U}$ as a ${\mathbb{Q}}$-divisor since there is no codimension 1 ramifications of $\pi$, and a line bundle $\pi^{*}E$ on $\widetilde{U}$. We consider
$\pi^{*}{\sigma}\in H^{0}(\widetilde{U},K_{\widetilde{U}/Y}+\pi^{*}E),$
which can be seen as a (usual) $\pi^{*}E$-valued relative holomorphic $n$-form on $\widetilde{U}$. We note that $(\pi^{*}{\sigma})^{p}=\pi^{*}({\sigma}^{p})$ holds for each positive integer $p$. The equality is valid over $\pi^{-1}(U_{reg})$ initially, and continue to hold on $\widetilde{U}$.
(2) Let $g$ be a singular Hermitian metric on the ${\mathbb{Q}}$-line bundle $E$ such that every local weight function $-\log g$ is quasi-psh. We define a (possibly singular) relative volume form
$\omega=g\,c_{n}{\sigma}\wedge\overline{\sigma}$
on $X$ as follows. In the first place, $\omega=gc_{n}{\sigma}\wedge\overline{\sigma}$ is a relative $(n,n)$-form on $X_{reg}$, where $c_{n}=\sqrt{-1}^{n^{2}}$. For each orbi-fold local chart $\pi:\widetilde{U}\to U=\widetilde{U}/G$, $\pi^{*}\omega$, defined on $\pi^{-1}(U_{reg})$, extends to $(\pi^{*}g)c_{n}\pi^{*}{\sigma}\wedge\overline{\pi^{*}{\sigma}}$ on $\widetilde{U}$, where $\pi^{*}g$ is a $G$-invariant singular Hermitian metric of $\pi^{*}E$ and $\pi^{*}{\sigma}$ is a $G$-invariant $\pi^{*}E$-valued relative (with respect to $f\circ\pi$) holomorphic $n$-form on $\widetilde{U}$. We still denote $\pi^{*}\omega=(\pi^{*}g)c_{n}\pi^{*}{\sigma}\wedge\overline{\pi^{*}{\sigma}}$ on $\widetilde{U}$.
(2’) Let $(V,t=(t_{1},\ldots,t_{m}))$ be a local coordinate of $Y$ and put $X_{V}=f^{-1}(V)$. We trivialize $K_{Y}$ on $V$ by $dt=dt_{1}\wedge\ldots\wedge dt_{m}$. We take a volume form $|dt|^{2}=c_{m}dt\wedge\overline{dt}$ on $V$. We then understand $\omega\wedge f^{*}|dt|^{2}$ as a possibly singular semi-positive $(n+m,n+m)$-form (a positive measure) on $X_{V}$. It is initially defined on $X_{V}\cap X_{reg}$ as a positive measure. On each orbi-fold local chart $\pi:\widetilde{U}\to U=\widetilde{U}/G$ such that $U\subset X_{V}$, we have $\pi^{*}{\sigma}\in H^{0}(\widetilde{U},K_{\widetilde{U}/Y}+\pi^{*}E)={\rm Hom}_{{\mathcal{O}}_{\widetilde{U}}}((f\circ\pi)^{*}K_{Y},K_{\widetilde{U}}+\pi^{*}E)$. Then we have $(\pi^{*}{\sigma})((f\circ\pi)^{*}dt)\in H^{0}(\widetilde{U},K_{\widetilde{U}}+\pi^{*}E)$, and have a positive measure $(\pi^{*}g)c_{n+m}(\pi^{*}{\sigma})((f\circ\pi)^{*}dt)\wedge\overline{(\pi^{*}{\sigma})((f\circ\pi)^{*}dt)}$ on $\widetilde{U}$ which is $G$-invariant. The pull-back $\pi^{*}(\omega\wedge f^{*}|dt|^{2})$ over $U_{reg}$ extends to $(\pi^{*}g)c_{n+m}(\pi^{*}{\sigma})((f\circ\pi)^{*}dt)\wedge\overline{(\pi^{*}{\sigma})((f\circ\pi)^{*}dt)}$.
Regarding $\omega\wedge f^{*}|dt|^{2}$ as an $(n+m,n+m)$-form on $X_{V}$, we can consider its tensor power $(\omega\wedge f^{*}|dt|^{2})^{p}$ for every $p\in{\mathbb{Z}}_{>0}$. We can not integrate it if $p>1$, but we can compare such powers of forms pointwise on $X_{V}$.
(3) Denote $Y_{f\,\text{q-sm}}=\{y\in Y;\ \text{$f$ is quasi-smooth over $y$}\}$. Here $f$ is quasi-smooth at $x\in X$, if the orbi-fold tangent map $df:T_{X,x}\to f^{*}T_{Y,f(x)}$ is surjective, equivalently $d(f\circ\pi):T_{\widetilde{U},\widetilde{x}}\to(f\circ\pi)^{*}T_{Y,f(x)}$ is surjective for a(ny) $\widetilde{x}\in\widetilde{U}$ such that $\pi(\widetilde{x})=x$, where $\pi:\widetilde{U}\to U/G$ is an orbi-fold local coordinate around $x$. We define a function $F_{\omega}:Y_{f\,\text{q-sm}}\to[0,+\infty]$ given by the fiberwise ($L^{2}$-)integral of ${\sigma}$ with respect to $g$:
$F_{\omega}(y)=\int_{X_{y}}\omega\ \in[0,+\infty].$
We take an open covering $\{U_{k}\}$ of $X$, which is locally finite and each $U_{k}$ is a quotient $\pi_{k}:\widetilde{U}_{k}\to U_{k}=\widetilde{U}_{k}/G_{k}$ as above. We take a partition of unity $\{\rho_{k}\}$ subordinate to $\{U_{k}\}$. We then define, for every $y\in Y_{f\,\text{q-sm}}$,
$F_{\omega}(y)$ $:=\int_{X_{y}}\omega=\sum_{k}\int_{X_{y}\cap U_{k}}\rho_{k}\omega,$
$\int_{X_{y}\cap U_{k}}\rho_{k}\omega$ $:=\frac{1}{|G_{k}|}\int_{(f\circ\pi_{k})^{-1}(y)}\pi_{k}^{*}(\rho_{k}\omega)\ \in[0,+\infty].$
(3’) Let $(V,t=(t_{1},\ldots,t_{m}))$ be a local coordinate of $Y$ and put $X_{V}=f^{-1}(V)$. We have an $(n+m,n+m)$-form $\omega\wedge f^{*}|dt|^{2}$ on $X_{V}$. Then the push-forward is
$f_{*}(\omega\wedge f^{*}|dt|^{2})=F_{\omega}|dt|^{2}$
on $V\cap Y_{f\,\text{q-sm}}$. ∎
In our main concern, we take those things as follows (in 3.1.2 and 3.2.3, for example). We will also use a version that is localized over the base $Y$ (in the proof of 4.2, for example).
3.2.2.
Let everything be as in 1.1. We then take objects $E,{\sigma},\ldots$ in 3.2.1 as follows.
(1) We take an orbi-line bundle
$E=\ulcorner{D_{-}}\urcorner-K_{X/Y}=\ulcorner{D_{-}}\urcorner-K_{X}+f^{*}K_{Y}.$
As $X$ is compact, $E$ is a ${\mathbb{Q}}$-line bundle. We see $({\Delta}_{D}+M)-E\sim_{\mathbb{Q}}f^{*}L$. We take a section
${\sigma}=s_{\ulcorner{D_{-}}\urcorner}\in H^{0}(X,{\mathcal{O}}_{X}(\ulcorner{D_{-}}\urcorner))=H^{0}(X,{\mathcal{O}}_{X}(K_{X/Y}+E))$
whose divisor is $\ulcorner{D_{-}}\urcorner=\ulcorner{-D}\urcorner$ (to be more precise the divisor of $s_{\ulcorner{D_{-}}\urcorner}|_{X_{reg}}$ is $\ulcorner{-D}\urcorner|_{X_{reg}}$). We take a (root of) section $s_{{\Delta}_{D}}$ whose divisor is ${\Delta}_{D}$, and any Hermitian metric $h_{{\Delta}_{D}}$ on the ${\mathbb{Q}}$-line bundle ${\Delta}_{D}$. Let $h_{M}$ be a Hermitian metric on the ${\mathbb{Q}}$-line bundle $M$, and ${\varphi}$ be a quasi-psh function on $X$. (A model case is that ${\varphi}$ has a vanishing Lelong number everywhere.) Let
$g=\frac{h_{{\Delta}_{D}}}{|s_{{\Delta}_{D}}|_{h_{{\Delta}_{D}}}^{2}}e^{-{\varphi}}h_{M}$
be a singular Hermitian metric of ${\Delta}_{D}+M$.
(2) We take a small local coordinate $V$ of $Y$. We may suppose $V$ is biholomorphic to a ball in $({\mathbb{C}}^{m},t=(t_{1},\ldots,t_{m}))$, and $V$ is a relatively compact subdomain in a larger domain on which all of our relevant objects are trivialized. Denote by $X_{V}=f^{-1}(V)$. We can suppose $L\sim_{{\mathbb{Q}}}0$ on $V$. We have
$E_{V}:=E|_{X_{V}}\sim_{{\mathbb{Q}}}({\Delta}_{D}+M)|_{X_{V}}$
on $X_{V}$. A singular Hermitian metric on $E_{V}$ induces a singular Hermitian metric on a ${\mathbb{Q}}$-line bundle $({\Delta}_{D}+M)|_{X_{V}}$ and vice versa. We give $E_{V}(\sim_{{\mathbb{Q}}}({\Delta}_{D}+M)|_{X_{V}})$ a singular Hermitian metric
$g_{V}=g|_{X_{V}}.$
We then have a possibly singular relative $(n,n)$-form $\omega_{V}=g_{V}c_{n}{\sigma}\wedge\overline{\sigma}$ on $X_{V}$, and a function $F_{V}:V\setminus Q\to[0,+\infty]$ given by
$F_{V}(y)=\int_{X_{y}}\omega_{V}.$
(3) We take a Hermitian metric $h_{E}$ on the ${\mathbb{Q}}$-line bundle $E$, and let
$g_{E}=h_{E}\ \frac{e^{-{\varphi}}}{|s_{{\Delta}_{D}}|_{h_{{\Delta}_{D}}}^{2}}$
be a singular Hermitian metric on $E$. Then we have a possibly singular relative $(n,n)$-form $\omega=g_{E}c_{n}{\sigma}\wedge\overline{\sigma}$ on $X$ and a function $F_{\omega}:Y\setminus Q\to[0,+\infty]$ given by
$F_{\omega}(y)=\int_{X_{y}}\omega.$
If we change a Hermitian metric $h_{E}$ on $E$ to another $\widetilde{h}_{E}$, we have another $\widetilde{F}_{\omega}:Y\to[0,+\infty]$. As we have $c^{-1}h_{E}\leq\widetilde{h}_{E}\leq ch_{E}$ as a Hermitian metric on $E$ for some constant $c\in{\mathbb{R}}_{\geq 1}$, we have $c^{-1}F_{\omega}\leq\widetilde{F}_{\omega}\leq cF_{\omega}$. In particular, the “singularities” of $F_{\omega}$ depend only on the zero divisors of ${\sigma}$ and of $s_{{\Delta}_{D}}$, and on the singularities of ${\varphi}$ and of the map $f$. The difference between (2) and (3) above is the choice of metrics $(E_{V},g_{V})$: semi-global/local over $Y$, or $(E,h_{E})$: global. The singularities of $F_{V}$ in (2) and $F_{\omega}|_{V}$ in (3) are “equivalent”.
The fiber integral $F_{\omega}(y)$ and $F_{V}(y)$ are computed on every orbi-fold local chart of $X$, and are reduced to a computation of fiber integrals on a smooth total space ($\widetilde{U}_{k}$ as in 3.2.1 (3)). Thus we have
Theorem 3.2.3.
Let everything be as in 1.1 (with quasi-smooth $X$) and suppose ${\Delta}_{0}=0$. Let ${\varphi}$ be a quasi-psh function on $X$ with a vanishing Lelong number everywhere on $X$. We follow the notations in 3.2.2 (3). Let $y_{0}\in Y$ and let $V\subset({\mathbb{C}}^{m},t=(t_{1},\ldots,t_{m}))$ be a local coordinate centered at a point $y_{0}\in Y$. Then
(1) Upper estimate. $F_{\omega}/|\prod_{j=1}^{m}t_{j}|^{c}\in L^{1}_{loc}(V,0)$ for any $c\in(0,2)$. In particular, if $-\log F_{\omega}$ is quasi-psh on a neighborhood of $y_{0}$, one has $\nu(-\log F_{\omega},y_{0})=0$.
(2) Lower estimate. There exist positive constants ${\delta}(<1)$ and $c_{1}$ such that, for any $t\in V\setminus Q$ with $|t|<{\delta}$, $F_{\omega}(t)\geq c_{1}$ holds.
We note that, for every orbi-fold local chart $\pi:\widetilde{U}\to U$ of $X$, $\pi^{*}{\varphi}$ is quasi-psh on $\widetilde{U}$ and $\nu(\pi^{*}{\varphi},\widetilde{x})=0$ for every $\widetilde{x}\in\widetilde{U}$, see 2.2.5.
3.3. Comparison of fiber integrals
We will compare the singularities of a fiber integral $F_{\omega}$ on the quasi-smooth total space $X$ in 3.2.2 (2) and the one on a smooth model $X^{\prime}$. Let ${\alpha}:X^{\prime}\to X$ be a resolution of singularities as in 2.1.4. We will study a certain fiber integral $F_{\omega^{\prime}}$ for $f^{\prime}=f\circ{\alpha}$, and will see the singularities of $F_{\omega}$ and $F_{\omega^{\prime}}$ are “equivalent”. Curvature properties of $M$ are irrelevant.
3.3.1.
Let everything be as in 1.1. We take a log-resolution ${\alpha}:X^{\prime}\to X$, and let $K_{X^{\prime}}+D^{\prime}={\alpha}^{*}(K_{X}+D)$ as in 2.1.4. We then take objects $E,{\sigma},\ldots$ in 3.2.1 for $f^{\prime}=f\circ{\alpha}:X^{\prime}\to Y$ as follows.
(1) We take
$E^{\prime}=\ulcorner{D^{\prime}_{-}}\urcorner-K_{X^{\prime}/Y},$
which is a line bundle on $X^{\prime}$. We have $({\Delta}_{D^{\prime}}+M^{\prime})-E^{\prime}\sim_{\mathbb{Q}}{f^{\prime}}^{*}L$. We note that $E^{\prime}\neq{\alpha}^{*}E$ in general, where $E$ is the orbi-line bundle in 3.2.2. We take a section
${\sigma}^{\prime}=s_{\ulcorner{D^{\prime}_{-}}\urcorner}\in H^{0}(X^{\prime},\ulcorner{D^{\prime}_{-}}\urcorner)=H^{0}(X^{\prime},K_{X^{\prime}/Y}+E^{\prime})$
whose divisor is $\ulcorner{D^{\prime}_{-}}\urcorner=\ulcorner{-D^{\prime}}\urcorner$. We take a (root of) section $s_{{\Delta}_{D^{\prime}}}$ whose divisor is ${\Delta}_{D^{\prime}}$, and take any Hermitian metric $h_{{\Delta}_{D^{\prime}}}$ on the ${\mathbb{Q}}$-line bundle ${\Delta}_{D^{\prime}}$. We take the pull-back metric $h_{M^{\prime}}={\alpha}^{*}h_{M}$ on the ${\mathbb{Q}}$-line bundle $M^{\prime}={\alpha}^{*}M$, and let ${\varphi}^{\prime}={\alpha}^{*}{\varphi}$ be a quasi-psh function on $X^{\prime}$, where $h_{M}$ and ${\varphi}$ are the one in 3.2.2 (1). Let
$g^{\prime}=\frac{h_{{\Delta}_{D^{\prime}}}}{|s_{{\Delta}_{D^{\prime}}}|_{h_{{\Delta}_{D^{\prime}}}}^{2}}e^{-{\varphi}^{\prime}}h_{M^{\prime}}$
be a singular Hermitian metric of ${\Delta}_{D^{\prime}}+M^{\prime}$. Note that $g^{\prime}\neq{\alpha}^{*}g$ in general.
(2) We take a small local coordinate $V$ of $Y$ as in 3.2.1 (2), and let $X^{\prime}_{V}={f^{\prime}}^{-1}(V)$. We have
$E^{\prime}_{V}:=E^{\prime}|_{X^{\prime}_{V}}\sim_{{\mathbb{Q}}}({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{V}}$
on $X^{\prime}_{V}$. We give $E^{\prime}_{V}(\sim_{{\mathbb{Q}}}({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{V}})$ a singular Hermitian metric
$g^{\prime}_{V}=g^{\prime}|_{X^{\prime}_{V}}.$
We have a possibly singular relative $(n,n)$-form $\omega^{\prime}_{V}=g^{\prime}_{V}c_{n}{\sigma}^{\prime}\wedge\overline{\sigma}^{\prime}$ on $X^{\prime}_{V}$. Denote $Y_{f^{\prime}\,\text{sm}}=\{y\in Y;\ \text{$f$ is smooth over $y$}\}$. We consider a function
$F^{\prime}_{V}:V\setminus Y_{f^{\prime}\,\text{sm}}\to[0,+\infty],\ \ F^{\prime}_{V}(y)=\int_{X^{\prime}_{y}}\omega^{\prime}_{V}.$
As we do not know if $f^{\prime}:X^{\prime}\to Y$ is well-prepared, we can not estimate $F^{\prime}_{V}$ directly as in 3.1.2.
Lemma 3.3.2.
Take a small local coordinate $(V,t=(t_{1},\ldots,t_{m}))$ of $Y$ as in 3.2.2 (2). We obtain fiber integrals $F_{V}$ as in 3.2.2 (1) and (2), and $F^{\prime}_{V}$ as in 3.3.1.
Then, there exists a constant $c\in{\mathbb{R}}_{\geq 1}$ such that $c^{-1}F_{V}\leq F^{\prime}_{V}\leq cF_{V}$ as a function on $(V\setminus Q)\cap Y_{f^{\prime}\,\text{\rm sm}}$. In particular, $F^{\prime}_{V}$ also satisfies 3.1.2 in an appropriate sense.
Proof.
We will see that our claim is exactly a consequence of the formula $K_{X^{\prime}}+D^{\prime}={\alpha}^{*}(K_{X}+D)$. It is important to note that the integrand $\omega^{\prime}_{V}$ of $F^{\prime}_{V}$ (to be more precise $\omega^{\prime}_{V}\wedge{f^{\prime}}^{*}|dt|^{2}$) is defined on $X^{\prime}_{V}={f^{\prime}}^{-1}(V)$, not only on $X^{\prime}_{V}\setminus Y_{f^{\prime}\,{\rm sm}}$.
We cover $X^{\prime}_{V}={f^{\prime}}^{-1}(V)$ by a finite number of local charts $U^{\prime}_{\ell}$ such that each ${\alpha}(U^{\prime}_{\ell})$ contained in an orbi-fold local chart of $X$ (as in 3.2.1). Let $U^{\prime}\subset X$ be one of $\{U^{\prime}_{\ell}\}$, and let $\pi:\widetilde{U}\to U=\widetilde{U}/G$ be an orbi-fold local chart of $X$ with ${\alpha}(U^{\prime})\subset U$. We denote $V_{0}=(V\setminus Q)\cap Y_{f^{\prime}\,\text{\rm sm}}$.
We will denote $\omega,\omega^{\prime},F_{\omega}$ and $F_{\omega^{\prime}}$ instead of $\omega_{V},\omega^{\prime}_{V},F_{V}$ and $F^{\prime}_{V}$ respectively. We compare $(m,m)$-forms $F_{\omega}|dt|^{2}$ and $F_{\omega^{\prime}}|dt|^{2}$ on $V_{0}$. We have a possibly singular $(n+m,n+m)$-form $\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2}$ on $X^{\prime}_{V}={f^{\prime}}^{-1}(V)$ (not only on ${f^{\prime}}^{-1}(V_{0}$)). We have $F_{\omega^{\prime}}|dt|^{2}=f^{\prime}_{*}(\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2})$ by the fiberwise integrations on which $f^{\prime}$ is smooth. Following the notations in Step 3 in the proof of 3.1.2 (where the total space is smooth), we have ${\sigma}^{\prime}({f^{\prime}}^{*}dt)|_{U^{\prime}}={\sigma}^{\prime}_{U^{\prime}}dx^{\prime}$ for some ${\sigma}^{\prime}_{U^{\prime}}\in H^{0}(U^{\prime},{\mathcal{O}}_{X^{\prime}})(\cong H^{0}(U^{\prime},E^{\prime}))$, where $E^{\prime}=\ulcorner{D^{\prime}_{-}}\urcorner-K_{X^{\prime}/Y}$ in 3.3.1 and where $dx^{\prime}$ is a local frame of $K_{X^{\prime}}|_{U^{\prime}}$. Then
$\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2}=A_{1}^{\prime}\frac{e^{-{\varphi}^{\prime}}}{|s_{{\Delta}_{D^{\prime}}}|^{2}}\big{|}{\sigma}^{\prime}({f^{\prime}}^{*}dt)\big{|}^{2}=A_{2}^{\prime}e^{-{\varphi}^{\prime}}\frac{|{\sigma}^{\prime}_{U^{\prime}}|^{2}}{|s_{{\Delta}_{D^{\prime}}}|^{2}}|dx^{\prime}|^{2}$
on $U^{\prime}$, where $A_{1}^{\prime},A_{2}^{\prime}\in C^{\infty}(\overline{U^{\prime}},{\mathbb{R}}_{>0})$ which reflect local trivializations. We have a similar expression for $F_{\omega}|dt|^{2}$ formally, namely
$\omega\wedge{f}^{*}|dt|^{2}=A_{1}e^{-{\varphi}}\frac{|{\sigma}_{U}|^{2}}{|s_{{\Delta}_{D}}|^{2}}|dx|^{2}$
on $U$, where $A_{1}\in C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$. However as $K_{X}$ is merely ${\mathbb{Q}}$-Cartier, we do not have a precise meaning of $dx$ (a local frame of $K_{X}$), nor ${\sigma}_{U}\in H^{0}(U,{\mathcal{O}}_{X})$. Things are justified after passing to a local orbi-fold covering or passing to a high power to make every divisor to be Cartier. As we need to compare with objects on $X^{\prime}$, we adopt the latter way. We take a large and divisible integer $p$, so that $pK_{X}$ is Cartier, for example, and take a local frame $k_{U,p}\in H^{0}(U,pK_{X})$. Then $|k_{U,p}|^{2/p}$ can be seen as a positive measure on $U$. To be more precise, we take a reference Hermitian metric $h_{p}$ on a line bundle ${\mathcal{O}}_{X}(pK_{X})$, and put $|k_{U,p}|^{2}=|k_{U,p}|_{h_{p}}^{2}\cdot h_{p}^{-1}$ (as usual). Alternatively, thinking $\pi^{*}k_{U,p}\in H^{0}(\widetilde{U},pK_{\widetilde{U}})$, $|\pi^{*}k_{U,p}|^{2/p}$ is a positive measure on $\widetilde{U}$ and is $G$-invariant, and hence it induces $|k_{U,p}|^{2/p}$ on $U$. Let us look at ${\sigma}^{p}\in H^{0}(X,pK_{X/Y}+pE)={\rm Hom}_{{\mathcal{O}}_{X}}(f^{*}pK_{Y},pK_{X}+pE)$. We have ${\sigma}^{p}(f^{*}dt^{\otimes p})|_{U}={\sigma}_{U,p}k_{U,p}$ for some ${\sigma}_{U,p}\in H^{0}(U,{\mathcal{O}}_{X})(\cong H^{0}(U,pE))$. Then
$(\omega\wedge{f}^{*}|dt|^{2})^{p}=A_{2}e^{-p{\varphi}}\frac{|{\sigma}_{U,p}|^{2}}{|s_{{\Delta}_{D}}|^{2p}}|k_{U,p}|^{2},$
where $A_{2}\in C^{\infty}(\overline{U},{\mathbb{R}}_{>0})$.
We can write ${\alpha}^{*}k_{U,p}=J_{{\alpha},p}(dx^{\prime})^{\otimes p}$ for some meromorphic function $J_{{\alpha},p}$ on $U^{\prime}$ whose divisor is $\mbox{{\rm div}}\,J_{{\alpha},p}=pK_{X^{\prime}/X}|_{U^{\prime}}$. (We need to take a power $p$ to formulate this relation.) In view of the formula of change of variables, we write as $|{\alpha}^{*}k_{U,p}|^{2/p}=|J_{{\alpha},p}|^{2/p}|dx^{\prime}|^{2}$. Let us look at the ratio on $U^{\prime}$:
$\frac{(\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2})^{p}}{{\alpha}^{*}(\omega\wedge{f}^{*}|dt|^{2})^{p}}=({A_{2}^{\prime}}^{p}/{\alpha}^{*}A_{2})\frac{|{\sigma}^{\prime}_{U^{\prime}}|^{2p}}{|s_{{\Delta}_{D^{\prime}}}|^{2p}}\cdot\frac{{\alpha}^{*}|s_{{\Delta}_{D}}|^{2p}}{{\alpha}^{*}|{\sigma}_{U,p}|^{2}}\cdot\frac{|dx^{\prime}|^{2p}}{|{\alpha}^{*}k_{U,p}|^{2}}.$
Note that the right hand side has no “zeros” or “poles”. In fact, the “divisor” of the right hand side is $-pD^{\prime}+{\alpha}^{*}pD-pK_{X^{\prime}/X}=0$, because of $K_{X^{\prime}}+D^{\prime}={\alpha}^{*}(K_{X}+D)$. Here, we used that $\mbox{{\rm div}}\,{\sigma}^{\prime}_{U^{\prime}}-\mbox{{\rm div}}\,s_{{\Delta}_{D^{\prime}}}=\ulcorner{D^{\prime}_{-}}\urcorner-(D^{\prime}_{+}+\ulcorner{D^{\prime}_{-}}\urcorner-D^{\prime}_{-})=-D^{\prime}$ on $U^{\prime}$, and similarly $\mbox{{\rm div}}\,{\sigma}_{U,p}-\mbox{{\rm div}}\,s_{{\Delta}_{D}}^{p}=-pD$ on $U$. Hence we have
$\frac{\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2})^{p}}{{\alpha}^{*}(\omega\wedge{f}^{*}|dt|^{2})^{p}}\in C^{\infty}(\overline{U^{\prime}},{\mathbb{R}}_{>0}),$
and $\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2}=A{\alpha}^{*}(\omega\wedge{f}^{*}|dt|^{2})$ for some $A\in C^{\infty}(\overline{U^{\prime}},{\mathbb{R}}_{>0})$ pointwise on each $U^{\prime}=U^{\prime}_{\ell}$. Thus we have $c\in{\mathbb{R}}_{\geq 1}$ such that
$c^{-1}{\alpha}^{*}(\omega\wedge{f}^{*}|dt|^{2})\leq\omega^{\prime}\wedge{f^{\prime}}^{*}|dt|^{2}\leq c{\alpha}^{*}(\omega\wedge{f}^{*}|dt|^{2})$
pointwise on $X^{\prime}_{V}={f^{\prime}}^{-1}(V)$. Then, by taking the push-forward $f^{\prime}_{*}=f_{*}\circ{\alpha}_{*}$ over $V_{0}$, namely taking fiberwise integrals, we have
$c^{-1}F_{\omega}|dt|^{2}\leq F_{\omega^{\prime}}|dt|^{2}\leq cF_{\omega}|dt|^{2}$
on $V_{0}=(V\setminus Q)\cap Y_{f^{\prime}\,\text{\rm sm}}$. ∎
4. Semi-positivity of the moduli part
We shall prove 1.1. We construct a “canonical” singular Hermitian metric on the ${\mathbb{Q}}$-line bundle $M_{Y}=L$. We adapt a relative Bergman kernel construction of adjoint type bundles, developed in [BP08], [BP10], [PT], [CH] for it. Theorem 1.1 can be stated more concretely as in 4.2 on a smooth model of $X$. We follow [CH, §3] for a general outline of the proof.
Set up 4.1.
We follow the construction in 2.1.4.
(1) Starting from the situation in 1.1, we take a log-resolution ${\alpha}:X^{\prime}\to X$ of the pair $(X,P)$, and put $f^{\prime}=f\circ{\alpha}:X^{\prime}\to Y$. We write as $K_{X^{\prime}}+D^{\prime}={\alpha}^{*}(K_{X}+D)$ for a ${\mathbb{Q}}$-divisor $D^{\prime}$ on $X^{\prime}$. The pair $(X^{\prime},D^{\prime})$ satisfies (1)–(4) in 2.1.4. In particular,
$K_{X^{\prime}}+D^{\prime}+M^{\prime}\sim_{{\mathbb{Q}}}{f^{\prime}}^{*}(K_{Y}+L),$
where $M^{\prime}={\alpha}^{*}M$. We set $D^{\prime}=\sum_{j}d^{\prime}_{j}P^{\prime}_{j}$, and consider the discriminant part ${\Delta}_{Y}^{\prime}$ and the moduli part $M_{Y}^{\prime}=L-{\Delta}_{Y}^{\prime}$ for $f^{\prime}$. By 2.1.4, ${\Delta}_{Y}^{\prime}={\Delta}_{Y}$ and $M_{Y}^{\prime}=M_{Y}$. As we can suppose that ${\Delta}_{Y}=0$, we have ${\Delta}_{Y}^{\prime}=0$ and $M_{Y}^{\prime}=L$.
(2) Let $\Sigma^{\prime}\subset Y$ be the minimum Zariski closed subset such that $Y\setminus\Sigma^{\prime}=\{y\in Y;\ f^{\prime}$ is smooth over $y$ and $h^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner)=1\}$.
(3) By our assumption, there exist a Hermitian metric $h_{M}$ on $M$ and a quasi-psh function ${\varphi}$ on $X$ such that a singular Hermitian metric $e^{-{\varphi}}h_{M}$ of $M$ has semi-positive curvature and that ${\varphi}$ has a vanishing Lelong number everywhere on $X$. Let $h_{M^{\prime}}={\alpha}^{*}h_{M}$ and ${\varphi}^{\prime}={\alpha}^{*}{\varphi}$. We write $D^{\prime}={\Delta}_{D^{\prime}}-\ulcorner{D^{\prime}_{-}}\urcorner$ as in 2.1.1. Let $h_{{\Delta}_{D^{\prime}}}$ be a Hermitian metric on ${\Delta}_{D^{\prime}}$, and $s_{{\Delta}_{D^{\prime}}}$ be a root of section whose divisor is ${\Delta}_{D^{\prime}}$. We let
$g^{\prime}=\frac{h_{{\Delta}_{D^{\prime}}}}{|s_{{\Delta}_{D^{\prime}}}|_{h_{{\Delta}_{D^{\prime}}}}^{2}}e^{-{\varphi}^{\prime}}h_{M^{\prime}}$
be the induced singular Hermitian metric on ${\Delta}_{D^{\prime}}+M^{\prime}$ with semi-positive curvature. ∎
Noting the relation $K_{X^{\prime}/Y}+D^{\prime}+M^{\prime}=K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}-\ulcorner{D^{\prime}_{-}}\urcorner$ in mind, we first consider an adjoint type ${\mathbb{Q}}$-line bundle $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$. We note that ${\mathcal{I}}(g^{\prime})={\mathcal{O}}_{X^{\prime}}$, by 2.2.7, as the pair $(X^{\prime},{\Delta}_{D^{\prime}})$ is klt and $\nu({\varphi}^{\prime},x^{\prime})=0$ for any $x^{\prime}\in X^{\prime}$.
Theorem 4.2.
Let everything be as in 4.1.
(1) The ${\mathbb{Q}}$-line bundle $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ admits a singular Hermitian metric $b^{-1}$ whose curvature current satisfies $\frac{\sqrt{-1}}{2\pi}\Theta_{b^{-1}}\geq[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]$, where $[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]$ is the current of integration over the effective cycle $\ulcorner{D^{\prime}_{-}}\urcorner$. More precisely, the restriction $b|_{X^{\prime}_{y}}$ is the Bergman kernel metric on $-(K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{y}}$ with respect to $({\Delta}_{D^{\prime}}+M^{\prime},g^{\prime})|_{X^{\prime}_{y}}$ for every $y\in Y\setminus\Sigma^{\prime}$.
(2) The ${\mathbb{Q}}$-line bundle ${f^{\prime}}^{*}L$ admits a singular Hermitian metric $\widetilde{h}$ with semi-positive curvature, which is canonically constructed via the isomorphism ${f^{\prime}}^{*}L\sim_{{\mathbb{Q}}}K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}-\ulcorner{D^{\prime}_{-}}\urcorner$ and $b^{-1}$. The corresponding singular Hermitian metric $h_{Y}$ on $L$ (such that ${f^{\prime}}^{*}h_{Y}=\widetilde{h}$) has semi-positive curvature.
[…]
(4) (We now use the assumption that ${\varphi}$ has a vanishing Lelong number everywhere on $X$.) In the setting of (3) above, one has $\nu(\psi_{V},y)=0$ for every $y\in V$. In particular, $L$ is nef by 2.2.2.
Proof.
Step 1. Preliminary discussion. We let, as in 3.3.1,
$E^{\prime}=\ulcorner{D^{\prime}_{-}}\urcorner-K_{X^{\prime}/Y},$
which is a line bundle on $X^{\prime}$. We have $K_{X^{\prime}/Y}+E^{\prime}=\ulcorner{D^{\prime}_{-}}\urcorner$ and $E^{\prime}\sim_{{\mathbb{Q}}}{\Delta}_{D^{\prime}}+M^{\prime}-{f^{\prime}}^{*}L$. In particular, we can regard $E^{\prime}$ to be pseudo-effective modulo ${f^{\prime}}^{*}L$ and $\ulcorner{D^{\prime}_{-}}\urcorner$ to be its adjoint line bundle. We take an auxiliary Hermitian metric $h_{K^{\prime}}$ of $K_{X^{\prime}/Y}$, and put
$h=h_{K^{\prime}}h_{{\Delta}_{D^{\prime}}}h_{M^{\prime}}$
a Hermitian metric on $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ with the curvature form
$\theta=\frac{\sqrt{-1}}{2\pi}\Theta_{h}\in c_{1}(K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}).$
The curvature current of $g^{\prime}$ on ${\Delta}_{D^{\prime}}+M^{\prime}$ is
$\frac{\sqrt{-1}}{2\pi}\Theta_{g^{\prime}}=[{\Delta}_{D^{\prime}}]+\frac{\sqrt{-1}}{2\pi}\Theta_{e^{-{\varphi}^{\prime}}h_{M^{\prime}}}\geq 0,$
where $[{\Delta}_{D^{\prime}}]$ stands for the current of integration over the effective ${\mathbb{Q}}$-cycle ${\Delta}_{D^{\prime}}$.
We will consider restrictions of those ${\mathbb{Q}}$-line bundles and metrics $g^{\prime}$ and $e^{-{\varphi}^{\prime}}$ to fibers. There are some fibers on which such restrictions do not behave well. There are proper Zariski closed subsets $\Sigma_{1},\Sigma_{2}\subset Y$ and pluri-polar subsets $\Sigma_{3},\Sigma_{4}\subset Y$ with the following properties (o)–(iv).
(o) $\Sigma_{1}\subset\Sigma_{2}\subset\Sigma_{3}\subset\Sigma_{4}\subset Y$. We actually have $\Sigma_{2}=\Sigma^{\prime}$ in the statement.
(i) Let $\Sigma_{1}=Y\setminus Y_{f^{\prime}\,\text{sm}}$. In particular, $f^{\prime}$ is smooth over $Y\setminus\Sigma_{1}$.
(ii) Let $\Sigma_{2}^{\prime}=\{y\in Y;\ h^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner)\neq 1={\rm rank}\,f_{*}{\mathcal{O}}_{X^{\prime}}(\ulcorner{D^{\prime}_{-}}\urcorner)\}$, and $\Sigma_{2}=\Sigma_{2}^{\prime}\cup\Sigma_{1}$. We may also let $\Sigma_{2}^{\prime}=\{y\in Y;\ h^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner)\neq 1,\text{${\mathcal{I}}(g^{\prime})$ is not flat over $y\in Y$}\}$. As ${\mathcal{I}}(g^{\prime})={\mathcal{O}}_{X^{\prime}}$, the second condition gives no restriction on $\Sigma_{2}$. This is just a reminder to apply the base change theorem.
(iii) Let $\Sigma_{3}^{\prime}=\{y\in Y\setminus\Sigma_{2};\ H^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner\otimes{\mathcal{I}}(g^{\prime}|_{X^{\prime}_{y}}))\subsetneqq H^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner\otimes{\mathcal{I}}(g^{\prime})|_{X^{\prime}_{y}})\}$, and $\Sigma_{3}=\Sigma_{3}^{\prime}\cup\Sigma_{2}$. We know ${\mathcal{I}}(g^{\prime})={\mathcal{O}}_{X^{\prime}}$, but we use the notation ${\mathcal{I}}(g^{\prime})$ to remind the restriction property of multiplier ideal sheaves. In particular, if $y\in Y\setminus\Sigma_{2}$ and $g^{\prime}|_{X^{\prime}_{y}}\equiv+\infty$, then $y\in\Sigma_{3}^{\prime}$. We have also $h^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner\otimes{\mathcal{I}}(g^{\prime}|_{X^{\prime}_{y}}))=0$ and $h^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner\otimes{\mathcal{I}}(g^{\prime})|_{X^{\prime}_{y}})=1$ for $y\in\Sigma_{3}^{\prime}$. The set $\Sigma_{3}$ is pluri-polar in $Y$ (see (iv) below).
(iv) We set $\Sigma_{4}^{\prime}:=\{y\in Y\setminus\Sigma_{2};\ {\mathcal{I}}(g^{\prime}|_{X^{\prime}_{y}})\subsetneqq{\mathcal{I}}(g^{\prime})|_{X^{\prime}_{y}}={\mathcal{O}}_{X^{\prime}_{y}}\}$, and $\Sigma_{4}=\Sigma_{4}^{\prime}\cup\Sigma_{2}$. This set $\Sigma_{4}$ is pluri-polar in $Y$ by [Xi, 2.9]. In particular, $\Sigma_{4}$ is a set of Lebesgue measure zero in $Y$ (this property has been observed in [MZ]). We see $\Sigma_{3}^{\prime}\subset\Sigma_{4}^{\prime}$. Hence $\Sigma_{3}$ is also pluri-polar in $Y$ (which is actually the content of [Xi, Step 2.6.4]). The role of $\Sigma_{4}$ is auxiliary.
Step 2. Localization over the base. Let $\{Y_{i}\}_{i\in I}$ be a locally finite open cover of $Y$ such that every $Y_{i}$ is small enough and biholomorphic to a ball in ${\mathbb{C}}^{m},m=\dim Y$. Denote by $X^{\prime}_{i}={f^{\prime}}^{-1}(Y_{i})$. We can suppose $L\sim_{{\mathbb{Q}}}0$ on every $Y_{i}$. Hence, on every $X^{\prime}_{i}$, we have
$E^{\prime}_{i}:=E^{\prime}|_{X^{\prime}_{i}}\sim_{{\mathbb{Q}}}({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{i}},$
and $K_{X^{\prime}/Y}+E^{\prime}_{i}\sim_{{\mathbb{Q}}}K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ on $X^{\prime}_{i}$. A singular Hermitian metric on $E^{\prime}_{i}$ (resp. $K_{X^{\prime}/Y}+E^{\prime}_{i}$) induces a singular Hermitian metric on a ${\mathbb{Q}}$-line bundle $({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{i}}$ (resp. $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ on $X^{\prime}_{i}$) and vice versa. Via $E^{\prime}_{i}\sim_{{\mathbb{Q}}}({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{i}}$, we can regard
$h_{i}:=h_{K^{\prime}}h_{{\Delta}_{D^{\prime}}}h_{M^{\prime}}|_{X^{\prime}_{i}}=h|_{X^{\prime}_{i}}$
as a Hermitian metric on $K_{X^{\prime}/Y}+E^{\prime}_{i}$ with $\frac{\sqrt{-1}}{2\pi}\Theta_{h_{i}}=\theta$ on $X^{\prime}_{i}$. We also give $E^{\prime}_{i}(\sim_{{\mathbb{Q}}}({\Delta}_{D^{\prime}}+M^{\prime})|_{X^{\prime}_{i}})$ a singular Hermitian metric
$g^{\prime}_{i}=g^{\prime}|_{X^{\prime}_{i}}.$
We have $\frac{\sqrt{-1}}{2\pi}\Theta_{g^{\prime}_{i}}=[{\Delta}_{D^{\prime}}|_{X^{\prime}_{i}}]+\frac{\sqrt{-1}}{2\pi}\Theta_{e^{-{\varphi}^{\prime}}h_{M^{\prime}}}\geq 0$ and ${\mathcal{I}}(g^{\prime}_{i})={\mathcal{O}}_{X^{\prime}_{i}}$.
We can see that $\{h_{i}/g^{\prime}_{i}\}_{i\in I}$ gives a global singular Hermitian metric on $K_{X^{\prime}/Y}$, i.e., $h_{i}/g^{\prime}_{i}=h_{j}/g^{\prime}_{j}$ on $X^{\prime}_{i}\cap X^{\prime}_{j}$. In fact, on each $X^{\prime}_{i}$,
$\frac{h_{i}}{g^{\prime}_{i}}=h_{K^{\prime}}h_{{\Delta}_{D^{\prime}}}h_{M^{\prime}}\Big{(}\frac{h_{{\Delta}_{D^{\prime}}}}{|s_{{\Delta}_{D^{\prime}}}|_{h_{{\Delta}_{D^{\prime}}}}^{2}}e^{-{\varphi}^{\prime}}h_{M^{\prime}}\Big{)}^{-1}=h_{K^{\prime}}|s_{{\Delta}_{D^{\prime}}}|_{h_{{\Delta}_{D^{\prime}}}}^{2}e^{{\varphi}^{\prime}}.$
Step 3. Fiberwise normalization. We take a section
${\sigma}^{\prime}=s_{\ulcorner{D^{\prime}_{-}}\urcorner}\in H^{0}(X^{\prime},\ulcorner{D^{\prime}_{-}}\urcorner)=H^{0}(X^{\prime},K_{X^{\prime}/Y}+E^{\prime})$
whose divisor is $\ulcorner{D^{\prime}_{-}}\urcorner$. By the base change theorem, we have $H^{0}(X^{\prime}_{y},K_{X^{\prime}/Y}+E^{\prime}_{i})={\mathbb{C}}{\sigma}^{\prime}_{y}$ with ${\sigma}^{\prime}_{y}:=s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{X^{\prime}_{y}}$ for $y\in Y_{i}\setminus\Sigma_{2}$. For every $y\in Y_{i}\setminus\Sigma_{2}$, we consider its $L^{2}$-norm with respect to $g^{\prime}_{i}$, namely the fiber integral in our context:
$\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}^{2}=F^{\prime}_{i}(y)=\int_{X^{\prime}_{y}}g^{\prime}_{i}c_{n}{\sigma}^{\prime}_{y}\wedge\overline{{\sigma}^{\prime}}_{y}\in(0,+\infty].$
This $\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}<+\infty$ if and only if ${\sigma}^{\prime}_{y}\in H^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner\otimes{\mathcal{I}}(g^{\prime}|_{X^{\prime}_{y}}))$. We can also see that, for $y\in Y_{i}\setminus\Sigma_{2}$, $\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}=+\infty$ if and only if $y\in\Sigma_{3}^{\prime}$.
For $y\in Y_{i}\setminus\Sigma_{3}$ (namely $\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}<+\infty$), there exists $\tau_{i,y}\in H^{0}(X^{\prime}_{y},\ulcorner{D^{\prime}_{-}}\urcorner)$ such that
$\int_{X^{\prime}_{y}}g^{\prime}_{i}c_{n}\tau_{i,y}\wedge\overline{\tau_{i,y}}=1.$
Such $\tau_{i,y}$ is unique up to a multiplication of a complex number of unit norm. In fact, we can take
$\tau_{i,y}:=\frac{{\sigma}^{\prime}_{y}\ }{\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}}\in H^{0}(X^{\prime}_{y},K_{X^{\prime}/Y}+E^{\prime}_{i})$
for $y\in Y_{i}\setminus\Sigma_{3}$. We take $\tau_{i,y}=0\in H^{0}(X^{\prime}_{y},K_{X^{\prime}/Y}+E^{\prime}_{i})$ when $y\in Y_{i}\cap(\Sigma_{3}\setminus\Sigma_{2})=Y_{i}\cap\Sigma_{3}^{\prime}$ (namely $\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}=+\infty$). Hence, the formulation $\tau_{i,y}:=\frac{{\sigma}^{\prime}_{y}\ }{\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}}$ is still valid formally for any $y\in Y_{i}\setminus\Sigma_{2}$.
Step 3’. Role of fiber integral. We have a relative $(n,n)$-from $\omega^{\prime}_{i}=g^{\prime}_{i}c_{n}{\sigma}^{\prime}\wedge\overline{{\sigma}^{\prime}}$ on $X^{\prime}_{i}$ and the fiber integral
$F^{\prime}_{i}(y)=\int_{X^{\prime}_{y}}\omega^{\prime}_{i}\in[0,+\infty]$
for $y\in Y_{i}\cap Y_{\text{$f^{\prime}$\,sm}}$. We note that the integrand $\omega^{\prime}_{i}$ is defined on $X^{\prime}_{i}$, not only on $X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$.
Step 4. Fiberwise Bergman kernel metric. We now apply [BP08] ([PT]) for $f:X^{\prime}_{i}\to Y_{i}$ and $K_{X^{\prime}/Y}+E^{\prime}_{i}$ with respect to $(E^{\prime}_{i},g^{\prime}_{i})$.
We let $\tau_{i}=\{\tau_{i,y}\}_{y\in Y_{i}\setminus\Sigma_{2}}$ be a family of fiberwise holomorphic sections parametrized by $Y_{i}\setminus\Sigma_{2}$. We define an ${\mathbb{R}}_{\geq 0}$-valued function ${\beta}_{i}$ on $X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$ by
${\beta}_{i}(x^{\prime}):=|\tau_{i}|_{h_{i}}^{2}(x^{\prime})=|\tau_{i,y}|_{h_{i}}^{2}(x^{\prime})$
for $x^{\prime}\in X^{\prime}_{y}$ with $y=f^{\prime}(x^{\prime})\in Y_{i}\setminus\Sigma_{2}$. This ${\beta}_{i}$ takes a value 0 exactly alone $(\mbox{{\rm Supp}}\,D^{\prime}_{-}\cup{f^{\prime}}^{-1}(\Sigma_{3}))\cap(X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2}))$. We let
$b_{i}={\beta}_{i}h_{i}^{-1}.$
On each fiber $X^{\prime}_{y}$ for $y\in Y_{i}\setminus\Sigma_{2}$ (even if $y\in\Sigma_{3}$), $b_{i}|_{X^{\prime}_{y}}$ is a singular Hermitian metric on $(K_{X^{\prime}_{y}}+E^{\prime}_{i}|_{X^{\prime}_{y}})^{*}$, which is called the Bergman kernel metric with respect to the Hermitian line bundle $(E^{\prime}_{i},g^{\prime}_{i})|_{X^{\prime}_{y}}$. We have $b_{i}|_{X^{\prime}_{y}}\equiv 0$ for $y\in Y_{i}\cap\Sigma_{3}^{\prime}$ by construction. We consider its dual
$b_{i}^{-1}=\frac{h_{i}}{|\tau_{i}|_{h_{i}}^{2}}=h_{i}\,e^{-\psi_{i}}\ \text{ with }\ \psi_{i}=\log|\tau_{i}|_{h_{i}}^{2}$
defined on $X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$. We note that $b_{i}$ does not depend on the choice of a reference metric $h_{i}$ but depends (only) on $g^{\prime}_{i}$.
Then by [BP08, 0.1] (see also [PT, 4.1.1]), on every local chart of $X^{\prime}_{i}$, a function $\log b_{i}(\sim\psi_{i})$ is psh on $X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$ and uniformly bounded from above around ${f^{\prime}}^{-1}(\Sigma_{2})$. (The last uniform upper-boundedness of $\psi_{i}$ also follows from 3.3.2 and 3.2.3 (2) in our setting. See Step 5 below.) Hence, by a Riemann type extension for psh functions $b_{i}^{-1}$ extends as a singular Hermitian metric of $K_{X^{\prime}/Y}+E^{\prime}_{i}$ on $X^{\prime}_{i}$ (we still denote it by $b_{i}^{-1}=h_{i}\,e^{-\psi_{i}}$) with semi-positive curvature
$\theta+dd^{c}\psi_{i}\geq 0$
on $X^{\prime}_{i}$, where $dd^{c}=\frac{\sqrt{-1}}{2\pi}{\partial}{\overline{\partial}}$. We regard $\psi_{i}$ as a quasi-psh function on $X^{\prime}_{i}$ from now on. Via $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}\sim_{{\mathbb{Q}}}K_{X^{\prime}/Y}+E^{\prime}_{i}$ on $X^{\prime}_{i}$, we regard $b_{i}^{-1}$ as a singular Hermitian metric of $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ on $X^{\prime}_{i}$.
Step 5. Singularities of the dual of the Bergman kernel metric. We note that a simple relation
$e^{-\psi_{i}}|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{i}}^{2}(x^{\prime})=\frac{F^{\prime}_{i}(f^{\prime}(x^{\prime}))}{|s_{\ulcorner{D^{\prime}_{-}}\urcorner}(x^{\prime})|_{h_{i}}^{2}}|s_{\ulcorner{D^{\prime}_{-}}\urcorner}(x^{\prime})|_{h_{i}}^{2}=F^{\prime}_{i}(f^{\prime}(x^{\prime}))$
holds for any $x^{\prime}\in X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$. In particular, thanks to the lower bound in 3.3.2 and 3.2.3 (2), a uniform upper bound
$\psi_{i}\leq\log|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{i}}^{2}+\log c_{1}^{-1}$
holds on $X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2}\cup Q)$ for a constant $c_{1}>0$. As both $\psi_{i}$ and $\log|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{i}}^{2}$ are quasi-psh on $X^{\prime}_{i}$, the relation $\psi_{i}\leq\log|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{i}}^{2}+\log c_{1}^{-1}$ still holds on $X^{\prime}_{i}$. Moreover, thanks to the estimate $\psi_{i}\leq\log|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{i}}^{2}+\log c_{1}^{-1}$, we see that the Lelong number of $\psi_{i}$ at each point is greater than or equal to that of the current $[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]$ (namely the multiplicity of the divisor). Thus we have in fact
$\frac{\sqrt{-1}}{2\pi}\Theta_{h_{i}e^{-\psi_{i}}}=\theta+dd^{c}\psi_{i}\geq[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]$
as currents on $X^{\prime}_{i}$.
We now claim that $\psi_{i}=\psi_{j}$ on $X^{\prime}_{i}\cap X^{\prime}_{j}$. It is shown in [CH, 3.4, Step 4] (from page 12, line 20 to page 13, line 2) that, for every $y\in Y_{i}\cap Y_{j}\setminus\Sigma_{3}$, $\psi_{i}|_{X^{\prime}_{y}}=\psi_{j}|_{X^{\prime}_{y}}$ holds.
(Our $(E^{\prime}_{i},g^{\prime}_{i})$ corresponds to $(L_{i},h_{L_{i}})$ in [CH, p.11, line 6 from the bottom]. $h_{i}$ is the same [CH, 3.4, Step 2]. The Zariski open set $U_{i,0}\subset U_{i}(\subset Y)$ in [CH, p.11, the bottom line] is our $Y_{i}\setminus\Sigma_{2}\subset Y_{i}$. $\tau_{i,y}$ corresponds to $s_{y}(=s_{y,i})$ in [CH, p.12, line 1]. $\psi_{i}$ correspond to $2{\varphi}_{i}$ in [CH, p.12, line 8].)
Thus we have $\psi_{i}=\psi_{j}$ on $(X^{\prime}_{i}\cap X^{\prime}_{j})\setminus{f^{\prime}}^{-1}(\Sigma_{3})$. As both $\psi_{i}$ and $\psi_{j}$ are quasi-psh on $X^{\prime}_{i}\cap X^{\prime}_{j}$, and as ${f^{\prime}}^{-1}(\Sigma_{3})$ is pluri-polar in $X^{\prime}$, in particular the Lebesgue measure zero, $\psi_{i}=\psi_{j}$ on $X^{\prime}_{i}\cap X^{\prime}_{j}$ ([Gu, §K, 15 Theorem, 16 Corollary]). When $y\in Y_{i}\cap Y_{j}\cap(\Sigma_{3}\setminus\Sigma_{2})$, we have in particular $\psi_{i}|_{X^{\prime}_{y}}\equiv-\infty\equiv\psi_{j}|_{X^{\prime}_{y}}$.
Thus $\{\psi_{i}\}_{i}$ defines a quasi-psh function on $X^{\prime}$, which we denote by
$\psi\in L^{1}_{\rm loc}(X^{\prime},{\mathbb{R}}).$
Step 6. Metric with semi-positive curvature on $M_{Y}=L$. We first have a global singular Hermitian metric $e^{-\psi}h_{K^{\prime}}h_{{\Delta}_{D^{\prime}}}h_{M^{\prime}}$ on $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}$ with the curvature
$\theta+dd^{c}\psi\geq[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]$
on $X^{\prime}$. For a line bundle ${\mathcal{O}}_{X^{\prime}}(\ulcorner{D^{\prime}_{-}}\urcorner)$, we have a canonical singular Hermitian metric $1/|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|^{2}=h_{\ulcorner{D^{\prime}_{-}}\urcorner}/|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|_{h_{\ulcorner{D^{\prime}_{-}}\urcorner}}^{2}$ for any Hermitian metric $h_{\ulcorner{D^{\prime}_{-}}\urcorner}$ of ${\mathcal{O}}_{X^{\prime}}(\ulcorner{D^{\prime}_{-}}\urcorner)$. Then we consider a singular Hermitian metric $\widetilde{h}$ on $K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}-\ulcorner{D^{\prime}_{-}}\urcorner\sim_{{\mathbb{Q}}}{f^{\prime}}^{*}L$ given by
$\widetilde{h}=e^{-\psi}h_{K^{\prime}}h_{{\Delta}_{D^{\prime}}}h_{M^{\prime}}|s_{\ulcorner{D^{\prime}_{-}}\urcorner}|^{2}.$
The curvature current is
$\frac{\sqrt{-1}}{2\pi}\Theta_{\widetilde{h}}=\theta+dd^{c}\psi-[\,\ulcorner{D^{\prime}_{-}}\urcorner\,]\geq 0$
on $X^{\prime}$. On each fiber $X^{\prime}_{y}$, we have $({f^{\prime}}^{*}L)|_{X^{\prime}_{y}}\sim_{{\mathbb{Q}}}{\mathcal{O}}_{X^{\prime}_{y}}$, and hence the semi-positively curved singular Hermitian metric on $({f^{\prime}}^{*}L)|_{X^{\prime}_{y}}$ needs to be constant (possibly +$\infty$+ ∞). Thus, we can find a singular Hermitian metric $h_{Y}$ on $L$ with semi-positive curvature such that ${f^{\prime}}^{*}h_{Y}=\widetilde{h}$. We usually write it as $h_{Y}=f^{\prime}_{*}\widetilde{h}$. At each point $x^{\prime}\in X^{\prime}_{i}\setminus{f^{\prime}}^{-1}(\Sigma_{2})$, we have
$\widetilde{h}(x^{\prime})=e^{-\psi_{i}(x^{\prime})}h_{i}(x^{\prime})|s_{\ulcorner{D^{\prime}_{-}}\urcorner}(x^{\prime})|^{2}=\frac{F^{\prime}_{i}(f^{\prime}(x^{\prime}))}{|s_{\ulcorner{D^{\prime}_{-}}\urcorner}(x^{\prime})|_{h_{i}}^{2}}h_{i}(x^{\prime})|s_{\ulcorner{D^{\prime}_{-}}\urcorner}(x^{\prime})|^{2}=F^{\prime}_{i}(f^{\prime}(x^{\prime})).$
Needless to say, the right hand side $F^{\prime}_{i}(f^{\prime}(x^{\prime}))$ is constant on each fiber (it can be $\equiv$+$\infty$ + ∞ on some fibers). Thus $h_{Y}(y)=F^{\prime}_{i}(y)(=\|{\sigma}^{\prime}_{y}\|_{g^{\prime}_{i}}^{2})$ for every $y\in Y_{i}\setminus\Sigma_{2}$ and every $i$.
Step 7. Bounding the singularities of the metric: the nefness of $M_{Y}=L$. Let us recall that ${\varphi}$ has a vanishing Lelong number everywhere on $X$. Then, 3.3.2 and 3.2.3 (1) imply that the Lelong number of the local weight functions of $h_{Y}:\nu(-\log h_{Y},y)=0$ for every $y\in Y$. As a conclusion, the singular Hermitian metric $h_{Y}=f^{\prime}_{*}\widetilde{h}$ on $M_{Y}=L$ is “nef” by 2.2.2. ∎
Remark 4.3.
There are two points we used 3.1.2 (or 3.2.3) in the proof of 4.2.
(1) 3.1.2 (2) is used in Step 5. This follows from the existence of a positive lower bound of $e^{-{\varphi}}$ and a condition ${\Delta}_{Y^{\prime}}=0$, which is a generic condition on the special fiber.
(2) 3.1.2 (1) is used in Step 7. We used ${\Delta}_{Y^{\prime}}=0$ and $\nu({\varphi}^{\prime},x^{\prime})=0$ for every $x^{\prime}\in X^{\prime}$. ∎
Remark 4.4.
Up to Step 6 in the proof above, our arguments (on the proof of 4.2 (1)–(3)) also work under a weaker assumption that the singular Hermitian metric $e^{-{\varphi}^{\prime}}h_{M^{\prime}}$ has semi-positive curvature and the multiplier ideal ${\mathcal{I}}(g^{\prime})={\mathcal{O}}_{X^{\prime}}$ over a non-empty Zariski open subset $Y_{M}$ of $Y$; ${\mathcal{I}}(g^{\prime})={\mathcal{O}}_{X^{\prime}}$ on ${f^{\prime}}^{-1}(Y_{M})$. ∎
5. On finite generation problem of log canonical rings for generalized pairs
In this section, we apply twisted Kawamata’s semi-positivity theorem (Theorem 1.1) for the finite generation problem for some versions of canonical rings.
Definition 5.1.
A generalized pair $(X,B+M)$ consists of
  • a normal projective variety $X$,
  • an effective $\mathbb{R}$-divisor $B$ on $X$, and
  • a $-\mathbb{R}$-Cartier b-divisor over $X$ represented by some projective birational morphism $\varphi:X^{\prime}\to X$ and a nef $\mathbb{R}$-Cartier divisor $M^{\prime}$ on $X^{\prime}$
such that $M=\varphi_{*}M^{\prime}$ and $K_{X}+B+M$ is $\mathbb{R}$-Cartier.
Definition 5.2.
Let $(X,B+M)$ be a generalized pair. Let $Y\rightarrow X$ be a projective birational morphism. We can write
$K_{Y}+B_{Y}+M_{Y}=\pi^{*}(K_{X}+B+M),$
for some divisor $B_{Y}$ on $Y$. The generalized log discrepancy of $(X,B+M)$ at a prime divisor $E\subset Y$, denoted by $a_{E}(X,B+M)$ is defined to be $1-{\rm coeff}_{E}(B_{Y})$.
A generalized pair $(X,B+M)$ is said to be generalized log canonical (or glc for short) if all its generalized log discrepancies are non-negative. A generalized pair $(X,B+M)$ is said to be generalized Kawamata log terminal (or gklt for short) if all its generalized log discrepancies are positive. In the previous definitions, if the analogous statement holds for $M=0$, then we drop the word “generalized”.
First, we see the following lemma by [BCHM].
Lemma 5.3.
Let $(X,B+M)$ be a projective generalized klt pair such that $K_{X}+B+M$ is big. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
Proof.
Note that by [FM, Proposition 5.1], it is enough to show the lemma for a sufficiently large and divisible $l$. Taking a log resolution, we may assume that $(X,B)$ is simple normal crossing klt pair and $M$ is nef. Since $K_{X}+B+M$ is big, we take an effective divisor $E$ and an ample divisor $A$ such that $l(K_{X}+B+M)\sim_{\mathbb{Z}}E+A$. Since $M+\frac{1}{l}A$ is ample, we can choose an effective $\mathbb{Q}$-divisor $A^{\prime}$ such that $(X,B+\frac{1}{l}(E+A^{\prime}))$ is klt and $A^{\prime}\sim_{\mathbb{Z}}lM+A$ for a sufficiently large and divisible positive integer $l$. Since
$2l(K_{X}+B+M)\sim_{\mathbb{Z}}l(K_{X}+B+\frac{1}{l}(E+A^{\prime})),$
it is enough to show the finite generation of
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+\frac{1}{l}(E+A^{\prime}))),$
which follows from [BCHM, Corollary 1.1.2]. ∎
Using twisted Kawamata’s semi-positivity theorem 1.1, we prove the following theorem:
Theorem 5.4.
Let $(X,B+M)$ be a projective generalized klt pair such that a $\mathbb{Q}$-line bundle $M^{\prime}$ on a smooth higher birational model $X^{\prime}\to X$ has a semi-positive singular Hermitian metric $h_{M^{\prime}}$ with a vanishing Lelong number for the potential at every point of $X^{\prime}$. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
Proof.
When $\kappa(K_{X}+B+M)=0$ or $\kappa(K_{X}+B_{M})<0$,
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))\simeq\mathbb{C}[t]\text{ or }\mathbb{C},$
respectively, where $\mathbb{C}[t]$ is the polynomial ring with one variable. Thus we may assume that $\kappa(X,l(K_{X}+B+M))>0$. We consider the Iitaka fibration $f:X\dashrightarrow Y$of $l(K_{X}+B+M)$ and take a higher birational model $X^{\prime}\to X$ induces a holomorphic morphism $f^{\prime}:X^{\prime}\to Y$ which is birational equivalent to $f$. Thus, by replacing, we may assume that $(X,B)$ is a simple normal crossing klt pair admitting the Iitaka fibration $f:X\to Y$ of $K_{X}+B+M$, and that $M$ has a semi-positive singular Hermitian metric with a vanishing Lelong number for the potential at every point of $X$. Now, we apply the weak semi-stable reduction theorem for $f$.
We have a birational morphisms $\varphi:X^{\prime}\to X$ and $\psi:Y^{\prime}\to Y$, and a well prepared birational model $f^{\prime}:(X^{\prime},B^{\prime})\to Y^{\prime}$ of $f$ such that $M^{\prime}=\varphi^{*}M$ and
$f^{*}(K_{X}+B+M)=K_{X^{\prime}}+B^{\prime}+M^{\prime}.$
We use the notation in twisted Kawamata’s semi-positivity theorem 1.1. Consider the divisor $\Delta_{Y^{\prime}}^{-}$. Since $B$ is effective, any components of $f^{\prime*}\Delta_{Y^{\prime}}^{-}$ are $\varphi$-exceptional divisors. Thus, $f^{\prime}$ is also the Iitaka fibarion of $K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}$. Therefore the finite generation of $R(K_{X}+B+M)$ is reduced to that of $R(K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}).$ And that is equivalent to the finite generation of $R(K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}})$. Now $(Y^{\prime},\Delta_{Y^{\prime}}^{+})$ is klt and $M_{Y^{\prime}}$ is nef by Theorem 1.1. Siince $K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}}$ is big, the desired finite generation follows from Lemma 5.3. ∎
The following is an interesting corollary for even smooth projective varieties.
Corollary 5.5.
Let $X$ be a smooth projective variety such that anti-canonical bundle $\omega_{X}^{-1}$ has a semipositive singular metrics $h$ with a vanishing Lelong number for the potential at every point of $X$. Then, the anti-canonical ring
$\bigoplus_{m\geq 0}H^{0}(X,\omega_{X}^{-m})$
is a finite generated $\mathbb{C}$-algebra.
In the last, we discuss the finite generation of generalized canonical rings for generalized klt pairs in a general situation. Indeed we prove that for lower dimensional cases as follows:
Theorem 5.6.
Let $(X,B+M)$ be a projective generalized klt pair such that $\mathrm{dim}\,X\leq 3$. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
Proof.
If $\kappa(X,l(K_{X}+B+M))\leq 0$, the finite generation is clear as the proof of Theorem 5.4. We may assume that $1\leq\kappa(K_{X}+B+M)\leq 2$ by Lemma 5.3. The case of $\kappa(K_{X}+B+M)=1$ is known since the finite generation is always true for a section ring of a Cartier divisor with Iitaka dimension one (cf. [W, Lemma 4.4]). When $\mathrm{dim}\,X=2$, we complete the proof by Lemma 5.3 again. Thus we may assume that $\mathrm{dim}\,X=3$ and $\kappa(K_{X}+B+M)=2$. As the proof of Theorem 5.4, by taking the Iitaka fibration and log resolution, we may assume that there exists a projective surjective morphism $f:X\to Y$ with connected fibers such that $(X,B+M)$ is generalized klt and $\mathrm{dim}\,Y=2$. For a general fiber $F$ of $f$, it holds that $\kappa((K_{X}+B+M)|_{F})=0$ since the property of Iitaka fibrations. Since $F$ is smooth projective curve, we see that $(K_{X}+B+M)|_{F}\sim_{\mathbb{Q}}0.$ Thus we can apply [FiS, Theorem 2.20].
Then there exists $B_{Y}$ and $M_{Y}$ such that $(Y,B_{Y}+M_{Y})$ is generalized klt and the finite generation for $K_{X}+B+M$ is equivalent for that for a big divisor $K_{Y}+B_{Y}+M_{Y}$ as the proof of Theorem 5.4. Since $K_{Y}+B_{Y}+M_{Y}$ is big, the desired finite generation follows from Lemma 5.3. ∎